### v2

# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

Abstract
We present a low-cost legged mobile manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning purely in simulation. This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions and a low-level policy for quadruped movement and limb control, 2) a progressive exploration and learning approach that leverages privileged task decomposition information to train the teacher policy for long-horizon tasks, which will guide an imitation-based student policy for efficient training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
   In contrast to previous approaches that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer. When fully trained in simulation, a single policy autonomously solves long-horizon tasks such as search, move, grasp, and drop-into, achieving nearly 80% success. This performance is comparable to that of expert human teleoperation on the same tasks but operates in a more efficient way, at 1.5 times the speed of human expert. The sim-to-real transfer is fluid across diverse indoor and outdoor scenes under varying lighting conditions. Finally, we discuss the key techniques that enable the entire pipeline, including efficient RL training and sim-to-real, to work effectively for legged mobile manipulation, and present their ablation results.
Figure 1: SLIM in Real. Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in ground types, backgrounds, distractors, and other environmental factors. These scenes are not available in simulation training. The annotations at the bottom are added for understanding task progress and are not part of the input to the system.
I Introduction
Until recently, robots have primarily been limited to static, highly repetitive tasks, often in industrial manufacturing. As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environments - commonly referred to as mobile manipulators. Such systems can typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
There are several promising methodologies for training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer. Each of them has its pros and cons.
Imitation learning (IL) benefits from learning directly from human experts teleoperating the hardware, and thus reasonable behaviors and some success in a restricted domain can be achieved even with a handful of trajectories. However, diverse behavior data is required to improve the policy’s robustness even for a single task. This makes IL costly, especially if we further want to cover all the combinatorial variations of scenes and tasks for robot ability scaling.
Reinforcement learning (RL) can potentially generate an infinite amount of training data by allowing the robot to explore and interact with the environment. Furthermore, the resulting policy is often more optimal than a human teleoperating policy. However, directly training an RL policy in the real world can be challenging due to hardware maintenance, reward specification, and safety issues. Data scale is also largely limited by the difficulty of deploying a robot fleet for collecting data in parallel. As a result, mastering long-horizon tasks in the real world still poses a significant challenge.
RL in simulation could address the scaling, generalization, and safety issues by running massively parallel simulation with randomized scenes and tasks. But to successfully deploy the trained policy on a real robot, the challenge lies in sim-to-real gaps. Therefore, we attempt to develop effective and general techniques to bridge the gaps. This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost quadruped based robot to solve long-horizon, intricate manipulation tasks. The complexity of accurate manipulation on legs, the variations of the physical environment, the scene and task setups, and the long task horizon are some of the factors that compound and increase the difficulty of RL training. On top of these factors, limited hardware reliability further exacerbates the sim-to-real gap, making this a daunting task.
We train an end-to-end model that is able to achieve long-horizon tasks consisting of searching, grasping, transporting and dropping primitive tasks (Figure 1). For robot mobility, we choose a quadruped instead of a wheeled base, as it has the potential to adapt to many different types of terrains. The choice of a quadruped base results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue. To mitigate this, we utilize a hierarchical teacher-student design, consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also a teacher-student training framework [6] for the high-level policy. To the best of our knowledge, this is the first demonstration of a low-cost real-world quadruped manipulator fully trained with RL in simulation, for solving long-horizon tasks.
The teacher policy has access to privileged information (e.g., structured and accurate object features) that is unavailable in the real world. To tackle the difficulties of long-horizon task learning, we also incorporate task decomposition as a form of privileged information, where boundaries between subtasks are defined. The student policy is then trained to distill the teacher policy while maximizing task rewards, but with access to only standard sensor data such as camera feed and proprioceptive state. We also condition the student visuomotor policy on language instructions, providing a natural way to the user for specifying the task to be executed by the robot.
We name our robotic system SLIM (Sim-to-Real Legged Instructive Manipulation). We reiterate that SLIM is a complete robotic system with fluid sim-to-real transfer. Unlike prior work [30, 53, 59], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning. SLIM relies on pure RL, and thus it does not require a single demonstration trajectory. Our contributions are as follows:
  1. 1.
    We develop a low-cost quadruped manipulation system SLIM. To our best knowledge, SLIM is the first end-to-end system for solving long-horizon real-world mobile manipulation tasks from RL alone.
  2. 2.
    To address the long-horizon RL exploration challenge, we propose a hierarchical framework for training SLIM’s high-level mobile manipulation policy and low-level quadruped controller, and adopt a teacher-student setup for high-level training efficiency.
  3. 3.
    We identify several techniques that are crucial for SLIM to achieve successful sim-to-real transfer with inexpensive hardware, as shown in real-world ablation studies.
  4. 4.
    We conduct an extensive set of experiments with real-world evaluations, achieving higher success rate and faster task completion than baselines. We also demonstrate robust task success across a wide range of indoor and outdoor environments in real-world.
II Related Work
II-A Learning-based Quadruped Locomotion
Traditionally, quadruped locomotion has been tackled through classical control methods designed to follow hand-tuned gaits on flat ground [17, 9], dynamic rigid platforms [24], discrete terrain [2], and rough terrain [12]. Though impressive, such methods require significant human engineering efforts and can be brittle to environmental changes.
More recently, there has been explosive progress in using learning-based approaches for achieving quadruped locomotion. In particular, sim-to-real reinforcement learning (RL) has arisen as a robust solution showcasing impressive feats such as parkour [61, 8, 4] and speeds of up to 3.9m/s [35]. Other works have also focused on exploiting the strengths of legged locomotion over their wheeled counterparts through utilizing the strong inherent coupling of proprioception with egocentric vision [16, 1] and active estimation of the environment [37, 28, 31]. Furthermore, minimal tuning is necessary compared to classical approaches [14] with massively parallelized simulations having been shown capable of learning gaits in just minutes [43]. Finally, data-driven legged locomotion has shown great promise in its ability to zero-shot generalize to new morphologies [13], surfaces [36, 28, 31] and agile skills [50, 27].
II-B Legged Mobile Manipulation
Significant advancements in robust locomotion have further enabled researchers to push the boundaries of legged mobile manipulation. Building upon data-driven approaches, quadrupeds have been demonstrated pushing objects with their body [26], dribble around balls [27], and manipulate objects using egocentric [49] or calve mounted grippers [29]. Still, the most popular setup has been the traditional top-mounted 6DOF manipulator design [15, 18, 30, 59, 53, 34, 38, 45]. In addition to significant workplace expansion, manipulators offer other nuanced benefits, such as assisting with balance [22] and serving as an intuitive interface for collecting human demonstrations [18]. Classical approaches to solving top-mounted legged mobile manipulation consist of designing wholebody controllers [45] or even combining locomotion learning with model-based manipulation control [34]. Others simplify the need for low-level controllers by using high-level API calls ($\dot{x},\dot{\theta}$) provided by their respective quadruped robots to achieve graspability-aware policies [59], navigational pick-and-place [53], and language-conditioned mobile manipulation [40]. Complex tasks have also been achieved by leveraging teacher-student training setups in simulation for grasping [30] and door opening [60]. Finally, to expand the robot’s workspace even further, wholebody loco-manipulation has arisen where leg joints are actuated in a way to assist manipulation [15, 30, 38].
Most similar to our work, Liu et al. [30] introduced VBC, which trains a sim-to-real visuomotor whole-body loco-manipulation policy using a high-low hierarchical model and teacher-student training setup for the object pickup task. They achieve impressive results across a diverse range of objects by leveraging whole-body manipulation, though this approach comes with several key caveats. First, VBC requires the user to manually click on the target object to perform an initial segmentation using a third-party general-purpose model. This segmentation is then used to track the object through another third-party model. Consequently, to ensure consistent segmentation, the object must always remain within the frame. This requirement inherently limits the feasible starting state distribution of VBC, as the target object must be placed within a short distance in front of the robot and in view. Finally, due to the close proximity and visibility of the object, the VBC task horizon is relatively short, ending as soon as the object is lifted. In comparison,
  1. 1.
    SLIM is a complete and self-contained system. All modules–vision, high-level visual loco-manipulation, low-level motor control–are trained within a single framework, resulting in both minimal latency and full autonomy.
  2. 2.
    SLIM is fluid and intuitive. Conditioned on language, SLIM is able to rapidly accomplish tasks in a wide variety of scenes, requiring just a user language command.
  3. 3.
    Finally, SLIM is practical. With SLIM, we demonstrate impressive real-world, long-horizon task completion for a multi-stage mobile search and manipulation task.
III Architecture Overview
[…]
Figure 2: Hierarchical Framework and Visuomotor Policy Pipeline (Left). Given a language instruction and various sensor information, the policy generates two control signals: 1) arm joint position targets and 2) a locomotion command. The first control signal is directly passed to the arm driver (passthrough), while the second signal is passed into a low-level locomotion policy to generate leg joint position targets. Low-Level and High-Level Teacher-Student Policy Training (Right). SLIM’s training pipeline can be decomposed into the following three sequential stages of policy training. First, a low-level locomotion policy is trained to track a specified locomotion command via RL. Next, a high-level teacher policy is trained with privileged information via RL to solve the long-horizon task. Finally, the student visuomotor policy is trained by distilling the teacher policy while maximizing task rewards, conditioned on visual inputs. Both the teacher and student policies use the same frozen locomotion policy from the first stage to execute their locomotion commands.
III-A Observations and Actions
Observations.
The robot observation at each time step consists of three components:
  • $\mathbf{l}$
    A tokenized instruction vector of length $L$.
  • $\mathbf{o}$
    A temporal stack of RGB images from a single wrist-mounted camera with a shape of $(N,H,W,3)$, where $N$ is the stack size, and $H$ and $W$ are the image height and width, respectively. The environment is always partially observed by the robot.
  • $\mathbf{s}_{\textrm{r}}$
    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive reading of all joints. $\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive states and $s_{\textrm{arm}}$ the arm-related proprioceptive states. When used as the input to the low-level locomotion policy, we overload $\mathbf{s}_{\textrm{leg}}$ to denote the leg-related proprioceptive states stacked according to the low-level frame-stacking size.
Privileged Observations (only available in simulation).
To achieve better learning efficiency, we will first train another policy using privileged observations, and then distill it to our robot policy conditioned on the standard observations. This guidance policy has a subtask id input $k$ in place of $\mathbf{l}$ and a lower-dimensional object observation $\mathbf{s}_{\textrm{p}}$ in place of $\mathbf{o}$:
  • $k$
    An integer in $[1,K]$ indicating which subtask the teacher is currently solving. $K$ is the maximal number of subtasks forming a long-horizon task.
  • $\mathbf{s}_{\textrm{p}}$
    A group of temporal object feature stacks, where each stack has a shape of $(N,Q_{m})$ and $0\leq m<M$ denotes the feature index. $M$ is the maximal number of task related objects in the scene. Note that object features are also partially observed and their visibility is always determined by the robot’s camera field of view.
The privileged information can be extracted or computed from the underlying simulator state.
Actions.
The robot policy outputs a triplet of actions:
  • $\mathbf{a}_{\textrm{arm}}$
    A vector in $[-z,z]^{6}$, indicating the delta changes to the arm joint positions. $z$ is the maximally allowed joint position change for a control interval.
  • $g$
    A target gripper position in $[0,1]$.
  • $\mathbf{a}_{\textrm{leg}}$
    The target quadruped joint positions in $\mathbb{R}^{12}$.
In total, a complete action $\mathbf{a}$ has 19 dimensions.
III-B Hierarchical Policy Structure
The system as a whole takes visual and instructional input and controls all joints of the quadruped and the arm, which is a daunting training task with high-dimensional input and output. We use a two-level hierarchical policy to divide the complexity of visual mobile manipulation training from legged locomotion training, similar to [30], illustrated in the left half of Fig. 2.
The high level takes visual, instructional, and proprioceptive inputs, and outputs a simple locomotion speed command and arm control command, while the low level takes proprioceptive input and outputs leg joint commands to track the locomotion command. The assumption is that quadruped locomotion control is largely independent of the high-level task semantics.
Intermediary Action.
To enable this two-level policy structure, we introduce an intermediary action as follows,
  • $\mathbf{c}$
    A vector in $\mathbb{R}^{2}$, containing the target forward and angular velocities for the quadruped.
With this, a high-level policy is defined as
$(\mathbf{l},\mathbf{o},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a}_{\text{hi}}\triangleq(\mathbf{a_{\textrm{arm}}},g,\mathbf{c})\in\mathbb{R}^{9},$
while the low-level policy is defined as
$(\mathbf{c},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{12}.$
Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, thus does not need to handle high dimensional image and language inputs.
Low-Level.
The low-level policy is a quadruped controller that generates joint position targets for PD control to follow a task-agnostic 2D command $\mathbf{c}$ [36]. The low-level policy is trained in simulation using PPO [44] and transferred to real. An illustration of the training can be found in the first block in the right column of Fig. 2. It is worth noting that we randomly sample both command $\mathbf{c}$ and arm joint actions $\mathbf{a_{\textrm{arm}}}$ during low-level training. Since the arm mounted on the quadruped can have varying poses during the task, the low-level command following ability must generalize to various arm configurations. That is, the quadruped base has to learn to keep balance while achieving locomotion commands regardless of the arm’s current joint positions and movements. Similar to most sim2real approaches, we also randomize the simulation environment during training. Besides widely-used domain randomization parameters as in [15], we further randomize the simulated delay of each sensor and foot softness of the robot, to better adapt to variations in real deployment. Rewards design and full details of the training setup can be found in Appendix A-B.
High-Level.
The high-level policy is responsible for outputting the command $\mathbf{c}$ and the arm control $(\mathbf{a_{\textrm{arm}}},g)$, conditioned on language $\mathbf{l}$ and the current stack of RGB images $\mathbf{o}$. As demonstrated in Fig. 2, the command $\mathbf{c}$ is forwarded to the low-level policy which then follows the received command for a number of time steps, before the high-level policy outputs the next command. For the manipulator, we opt to operate in joint space (as opposed to task space) and thus no inverse kinematics is needed. Once the low-level policy is trained, it is frozen and used as a base controller by the high-level policy. In the remainder of this paper, we will only talk about the high-level policy unless otherwise stated.
III-C Teacher-Student Training
Given the challenges of an enormous exploration space implied by long-horizon tasks, we adopt a teacher-student learning framework  [11, 30, 60] for better high-level training efficiency. The overall structure of the framework is shown in the second and third blocks of the right side of Fig. 2. As shown, the teacher is trained purely with RL from privileged, structured, and low-dimensional environment data $(k,\mathbf{s_{\textrm{p}}},\mathbf{s_{\textrm{r}}})$ that can only be obtained from the simulator. After a high performing teacher policy is learned, it is frozen and used to guide the student’s policy learning via a distillation loss. As the student will eventually be deployed in the real world, its inputs are consistent with what our robot system has. Besides policy distillation, we also use an RL objective to allow the student’s behavior being shaped by the same set of task rewards used by the teacher. Later, we will show improvements of this distillation-guided RL setting, compared to pure policy distillation (Section VIII-B Distillation Only baseline.). In the next two sections, we will provide more technical details of the teacher and student.
IV The Teacher: Progressive Policy Expansion for Long-Horizon Task Learning
IV-A Long-Horizon Task Learning Challenges
Since we are targeting at solving long-horizon tasks via RL, we need an effective approach to train a policy to solve long-horizon tasks successfully. There are are few key challenges that must be addressed for solving the long horizon tasks:
  1. 1.
    Continual exploration: When learning a long-horizon task via RL, there may be a number of intermediate milestones that must be sequentially achieved. Therefore, the teacher must be able to keep exploring at its frontier despite solving an intermediate milestone stage, in pursuit of solving the full task. Without the ability of continual exploration, it is more likely to converge to a sub-optimal solution that solves the task only up to some intermediate milestones, therefore cannot solve the full task completely.
  2. 2.
    Loss of plasticity and catastrophic forgetting: As the teacher progresses into a task, it needs to properly cope with both the loss of capacity/plasticity [32, 33] of the network due to the evolving nature of the problem as well as the data imbalance during long-horizon RL task exploration and training, without impacting skills that have already been acquired (catastrophic forgetting).
IV-B Progressive Policy Expansion
Given the challenges listed above, we build upon the idea of Policy Expansion (PEX) [58] originally used in the context of offline-to-online RL. Intuitively, PEX divides the full learning process into stages (e.g. offline and online stages in [58]) and uses dedicated networks for learning within each stage. This way, the trainer allocates new network capacity for continual exploration and learning of new skills without affecting any existing skill (catastrophic forgetting).
Here we generalize this idea for addressing the challenges in long-horizon task learning as mentioned in Section IV-A, by extending it to multiple stages. More concretely, we partition the full long horizon task into a sequence of stages (subtasks), and then perform a policy expansion operation for each new encountered subtask. By doing this, it can effectively combat the challenges in long-horizon task learning, keep exploring beyond the currently solved subtasks to solve the subsequent subtasks, circumvent the loss of plasticity by leveraging the newly allocated policy network  [^1] We use the term policy network to represent all the networks that are required for learning a policy. For example, in the context of Actor-Critic formulation, it encompasses both the actor and critic networks. for learning, and avoid catastrophic forgetting by retaining the policy networks for previous subtasks. Given these progressive multi-stage expansions, we dub this strategy Progressive Policy Expansion.
Formally, we represent the full long-horizon task $\mathcal{T}$ as a composition of a set of subtasks
$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})$ (1)
where $\mathcal{C}$ is a composition operator, assembling the set of subtasks $\{\tau^{k}\}$ into the full task $\mathcal{T}$. $K$ denotes the maximum number of subtasks. We also define a task decomposional operator as follows, mapping an input state to the corresponding subtask index:
$k=\mathcal{D}(\mathbf{s})\quad k\in\{1,2,\cdots,K\},$ (2)
where $\mathbf{s}=[\mathbf{s}_{\textrm{r}},\mathbf{s}_{\textrm{p}}]$. We use $\pi_{\rm tea}^{k}(\mathbf{a}|\mathbf{s})$ to denote the teacher sub-network that is responsible for solving its corresponding subtask $\tau^{k}$. When there is no confusion, we omit the subscript of the teacher network $\pi_{\rm tea}^{k}$ as $\pi^{k}$. Figure 3 provides graphical illustrations of the teacher sub-network structures. Each sub-network (Figure 3 left) takes low-dimensional proprioceptive and privileged observation as input and is responsible for learning to solve the corresponding subtask.
[…]
Figure 3: Teacher Policy Network Structure. On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP. In the case of actor-critic learning, both critic and actor networks have the same architecture, with the difference in that the actor network outputs an action distribution while the critic outputs a scalar value. On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$. Given an input id $k$, the $k$-th component network is activated in the computation, i.e. $\Pi[k]\equiv\pi^{k}$.
We can then solve the long horizon task $\mathcal{T}$ via Progressive PEX as follows. We initiate the exploration and learning with a single policy network $\Pi=\{\pi^{1}\}$. This policy network is responsible for learning to solve the initial subtask. Whenever a new subtask is encountered, a policy expansion operation is conducted, adding a new policy into the policy set $\Pi=\{\pi^{1},\pi^{2}\}$. By doing this progressively for the full task, we get
$\Pi$$\triangleq\{\pi^{k}\}_{k=1}^{K},$ (3)
as illustrated in Figure 3 right. A forward pass through $\Pi$ is carried out via the proper sub-policy by indexing the policy set with the task decompositional operator $\mathcal{D}$ conditioned on the current state
$\Pi(\mathbf{a}_{\text{hi}}|\mathcal{D},\mathbf{s})$ $=\pi^{\mathcal{D}(\mathbf{s})}(\mathbf{a}_{\text{hi}}|\mathbf{s}).$ (4)
There are many possible ways to implement the task decompositional operator $\mathcal{D}$. In our case, since it is used in teacher policy learning, we can leverage the fact that privileged observation can be used by the teacher policy and implement it as part of the privileged observation provided by the simulator, e.g., providing the subtask index $k$ directly in the privileged observation. Based this implementation, the training of $\Pi$ is closely related to multi-task RL [56, 46], where we sample a batch of multi-task-transitions (transitions containing the task index $k$) from the replay buffer and use each transition in the batch for training the sub-network associated with its corresponding task index. In this work, we train the teacher policy using a multi-task variant of SAC algorithm [19, 56, 46],
V The Student: Policy Distillation Guided RL
The student needs to perceive the surrounding environment through RGB images. There is no input regarding which subtask the student is currently solving, and it has to infer this information from the inputs. Because of this, the student learns only a single policy across the entire long-horizon task (no sub-task dedicated policy).
To tackle these challenges, we train the student by distilling the teacher’s multiple sub-task policies into a single task policy $\pi_{\rm stu}$. However, distillation will only give us a policy that works fine but not quite well. It could lead to the state drifting or out-of-distribution (OOD) issues when the policy is rolled out for a long horizon. There is also always an observation gap between the teacher and the student, which sometimes creates a difficulty in minimizing the distillation loss. Moreover, due to the differences in model architectures and inductive biases, the student might need a different way of mastering certain skills. Finally, distillation makes the student’s performance upper bounded by the teacher’s. Given these considerations, we choose to boost the student’s policy with RL under the same set of task rewards used by the teacher (Fig. 4).
We train the student by modifying SAC [19] to incorporate the distillation loss properly. First, we use a mixed rollout strategy to generate replay data. When rolling out a new episode, with a probability of $\beta$ we will sample from the student policy, or otherwise from the teacher policy. On one hand, we wish to use the teacher to efficiently generate high-performing transitions along the long horizon. On the other hand, we want to keep exploring with the student’s own policy. Second, following SACLite [55] we remove the entropy reward from policy evaluation. Then for policy improvement, we replace the entropy term with the distillation loss, and assign a fixed weight $\alpha$ to it:
$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$ (5)
To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std. for Gaussian). This makes our distillation loss the best of two worlds: imitation learning and entropy regularization.
[…]
Figure 4: A close look at the student’s architecture. Red dashed lines represent data flows that only exist for training in simulation. Solid lines denote the inference process of the student policy. For deployment, the three inputs are replaced with real-world observations.
When designing the student’s representation model, we specifically take vision sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning. While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the vision sim2real gap (Fig. 4). The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space. This potentially leads to a better generalization ability. While there might be many choices for the vision bottleneck, from discrete latent codes [20] to canonical images [25], we choose to use a pair of segmentation and depth maps. There are mainly two reasons for this choice. First, we believe that they contain the minimal visual information required to achieve our task, while without incurring much task information loss. Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a vision bottleneck. Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs. Note that only the robot and task-related objects appearing in the instruction will be predicted in the segmentation mask. All other irrelevant object pixels will be classified as “background”. After this, the segmentation and depth maps will be fed to the downstream policy for generating actions. Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [30, 53].
VI Task Implementation In Simulation
Figure 5: Task Decomposition as Privileged Information. Given a task, we decompose it into a sequential series of primitive subtasks. As shown above, the task of finding an object A, grasping it, and then dropping into an object B can be decomposed into a series of subtasks. In simulation, the subtask boundaries and transitions are well-defined, serving as a form of privileged information to the teacher. The teacher can leverage this information, and use a separate policy to learn each subtask, thereby conquering the long horizon task. This expertise is then distilled into the student, which is conditioned on both visual camera feed and language commands.
We use a long-horizon pick-and-place task as a concrete example in this work (c.f. Figure 2 left for an illustration of the task), and we will use this task as a concrete example for illustrating the task decomposition as well. Given a language instruction (e.g., “Drop the blue cube into the green basket.”), the robot needs to locate the target object, picking it up and then drop it into the target basket.
As mentioned previously in Section IV, SLIM employs a task decompositional approach paired with progressive PEX for training a teacher policy to solve the long horizon task. More concretely, for the long-horizon task under consideration, it can be naturally broken down into the following sequential steps (c.f. Figure 5):
  1. 1.
    Search: the robot should first search for the target object (the green cube) for picking up (a subtask), based on the RGB images from the gripper mounted ego-centric camera;
  2. 2.
    MoveTo: after finding the target object, the robot moves to the object until it is within the reach of the arm;
  3. 3.
    Grasp: then the robot moves the gripper towards the target object for picking it up. During the grasping process, the robot base should ideally move less compared to the MoveTo subtask to facilitate accurate grasping;
  4. 4.
    SearchWObj: after picking up the target object, the robot should keep it in the gripper and then search for the target basket; within this stage, there should be some coordination between arm and locomotion movement to facilitate search while avoid the cube to be dropped.
  5. 5.
    MoveToWObj: after finding the target basket, the robot should move towards it while keeping the object in gripper, until the target basket is within a range that is appropriate for the robot to reach for dropping the object in its gripper;
  6. 6.
    MoveGripperToWObj: the robot can now move the gripper towards the target basket until reaching an appropriate position on top of the basket for dropping;
  7. 7.
    DropInto: the robot releases its gripper and drops the cube into the basket.
In addition, we can further append an auxiliary Idle subtask, which encourage the robot to adjust its arm back to neutral position after dropping the object into the basket, facilitating subsequent deployment. This is an optional subtask that is not required for task completion. We include it in our task implementation to make the final policy behavior safe and more predictable after completing the full task. When evaluating the real world performance (e.g. Section VIII-D), we do not include this auxiliary subtask into the full task metric. We also exclude it whenever we discuss task-relevant decompositions since it is not a task related subtask.
Formally, we instantiate the task compositional operator as
$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})\triangleq\tau^{1}\rightarrow\tau^{2}\rightarrow\cdots\rightarrow\tau^{K}$ (6)
where $\tau^{i}\!\rightarrow\!\tau^{j}$ denotes that the two subtasks $\tau^{j}$ and $\tau^{j}$ are two adjacent subtasks and $\tau^{j}$ follows subtask $\tau^{i}$ in a sequential manner. In our case, $K\!=\!7$ and $\{\tau^{1},\tau^{2},\cdots,\tau^{7}\}$ corresponds to {Search, MoveTo, Grasp, SearchWObj, MoveToWObj, MoveGripperToWObj, DropInto}.
Because of the decomposition, we can implement reward functions for each subtask, making it modular and easier for implementation. From some tasks, we simply use the sparse subtask success reward (e.g. Search). We incorporated delta-distance based shaping reward for tasks that requires to move the robot (or gripper) close to a target region (e.g. MoveTo, MoveToWObj, MoveGripperToWObj, Grasp, DropInto), in addition to the sparse subtask reward. For SearchWObj and MoveToWObj, we also incorporated an arm retract reward, which is another instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f. Figure A.13). Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position. Its impact will be investigated in Section VIII-B (No Arm Retract baseline).
An additional benefit of task decomposition is that we have the option to incorporate behavior priors in a flexible way into the policy. More details on the subtasks, rewards and behavior priors are provided in Appendix A-A.
VII Sim-to-Real Gap Reduction Techniques
sim-to-real is notoriously difficult. This work handles two kinds of sim-to-real gaps: the physical dynamics and the visual gaps. For brevity, we discuss a subset of the techniques which are crucial in addressing sim-to-real and for successful real-world deployment. More techniques can be found in Appendix D.
VII-A Dynamics Gap Reduction
The dynamics gap is essentially caused by misaligned transition functions $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ between the simulated and real worlds. This difference in dynamics will result in the sim-trained policy experiencing covariate shift when deployed in the real world, which if large enough, can destroy policy performance. In fact, even minor covariate shifts can result in task failure if the task requires fine motor control, e.g., grasping a small object. Therefore, to minimize the dynamics gap, we must minimize the impact of the covariate shift when transferring to real. In particular, we use the following key techniques for addressing the issue: 1) arm joint tracking error minimization, 2) domain randomization, and 3) object perturbations.
Minimizing Arm Tracking Error via PID Control.
We used PID controller for the robot arm. As opposed to PD controllers which will always have gravity-induced tracking errors, the integral term of PID controllers allow for more robust pose-agnostic tracking given that the joint delta targets are sufficiently small for particular operating frequency. This allows us to greatly reduce the dynamics sim-to-real gap while avoiding system identification. More details can be found in Appendix D-A.
Arm Control Perturbation.
Random structured noise is added to the arm control signal to make the learned policy to be robust w.r.t. different sources of control noises arising when transferring to real.
Arm Mount Perturbation.
We randomly perturb the arm mount position and yaw to make the learned policy robust to the sim-to-realgaps that could arise due to the actual mounting accuracy, torso height variations etc..
Object Perturbations.
The final strategy we employ are object perturbations. Due to the relatively deterministic nature of simulations (in contrast to the high-entropy real world), we observed that policies often converge to deterministic strategies. For example, for picking up an object, a policy may learn to always position itself so that the object has the same relative position to its base. This deterministic “memorizing” behavior can result in task failure if the conditions for success are even slightly off. Therefore, we enforce learning robust, reactive policies by perturbing objects of interest during various stages of task execution. Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robust ness of the policy (as the state coverage is enlarged).
Stationary Manipulations.
We make the quadruped stand as still as possible when manipulating objects, i.e. during grasp or drop into. This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed. To achieve this, when training low level policies to follow locomotion speed commands, we also make the quadruped enter a stationary mode with a balanced stance when the speed command is close to zero. Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers and arm motors.
VII-B Visual Gap Reduction
The second source of sim-to-real gap for image-conditioned policies is the vision gap. This gap is caused by an RGB distribution mismatch between simulated and real world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios. Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
Image Domain Randomization.
Inspired by SECANT [11], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim-to-real transfer. Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image. We choose not to introduce higher-level image randomization/generation [54] because it might impact image semantics.
Random Textures.
We spawn the robot on an empty ground and randomize the ground texture. The texture can be arbitrary and generally have no specific semantics, as long as they provide enough visual distraction to the student. This distraction will force the student to learn to focus on the most important visual features that are related to the task at hand: objects with reasonable sizes and target colors. When deployed in real, we hope that this visual focus will get transferred and help the robot locate objects.
Random Spatial Augmentation.
To make the visual representation module robust, we incorporate spatial augmentations in the visual representation learning [51, 52].
Random Background Objects.
To further increase the visual complexity of the surrounding environment, we randomly sample and spawn a set of background objects in the room around the robot’s initial position. There are two categories of background objects: 1) primitive shapes (cuboid, cylinder, and sphere), and 2) Google scanned objects [10]. We make sure that these objects only serve the purpose of distraction and the student will not confuse them with task objects. It could collide with an background object during training with some penalty reward. When spawning an object, we also randomize its size, orientation, and texture (only for primitive shapes).
Refer to caption Refer to caption
Figure 6: An example scene populated with random background objects. Left: close-up view; Right: birds-eye view.
VIII Physical System, Experiments and Results
In this section, we briefly go over the robot system and the task used to evaluate SLIM.
VIII-A Physical Robotic System
For our robotic system, we use a Unitree Go1 with a top-mounted WidowX250S manipulator. An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount, as shown in Fig. 7, serving as the robot’s sole RGB vision feed. The depth stream is not used in this work. Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach. The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction. Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around \$12K USD. With the non-edu version, total cost is even lower ($\sim$6K USD) and reflects a bit more closely the cost of mass produced parts..
Figure 7: Qualitative Overview of the Robot System.
[…]
VIII-B Baselines
We compare SLIM with a number of baselines below:
  • No Arm Retract: removing the arm retract (raise) reward in RL training, which is used to encourage the robot’s arm to raise up after grasping the object (introduced in Section VI);
  • No Perturb: removing arm control randomization, arm mount perturbation and object perturbation (introduced in Section VII-A)
  • No Visual Aug: no background objects and no random spatial augmentation based visual augmentation (introduced in Section VII-B)
  • Distillation Only: agent training without using the RL loss (introduced in Section V), keeping only the distillation and representation loss.
  • Human Teleop: an expert human trained for half an hour in teleoperating the robot with locomotion and arm end-effector commands observing only the view of the wrist camera. While not a baseline per se, human teleoperation provides a valuable reference point. See Appendix B-D for more details.
We also tried another baseline No Distillation, which trains the visuomotor policy directly via RL without distillation from the teacher policy (i.e. RL for visuomotor policy learning from scratch). This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration. It is excluded from the tables.
VIII-C Evaluation Protocol
We use the Standard object spatial layout as shown in Figure 8 in a real-world scene Lobby. By doing so, it is easier to maintain repeatability across all methods and seeds. Besides the Standard in lobby, we also evaluate the robot in several other scenes with other layout. For task objects, we use cubes and baskets of different colors. More details on these physical objects are detailed in Appendix F-A.
For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene setup (i.e. different positioning and colors of objects) according to the evaluation protocol. For the detailed protocol, please refer to Appendix F-B.
TABLE I: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
Method Fully Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
Autonomous Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
[…]
No Visual Aug 98.3±2.498.3.2.498.3$\pm$ 2.498.3±2.4 88.3±4.788.3.4.788.3$\pm$ 4.788.3±4.7 63.3±26.663.3.26.663.3$\pm$ 26.663.3±26.6 56.6±22.556.6.22.556.6$\pm$ 22.556.6±22.5 52.9±18.752.9.18.752.9$\pm$ 18.752.9±18.7
Distillation Only 93.3±4.793.3.4.793.3$\pm$ 4.793.3±4.7 71.1±33.271.1.33.271.1$\pm$ 33.271.1±33.2 60.0±38.960.0.38.960.0$\pm$ 38.960.0±38.9 50.0±35.550.0.35.550.0$\pm$ 35.550.0±35.5 61.2±25.561.2.25.561.2$\pm$ 25.561.2±25.5
[…]
VIII-D Results and Analysis
[…]
Figure 8: Scene Spatial Layouts. Standard: we put the robot in the center of a square space and objects including objects to be grasped and containers to drop the graspable objects into on the four corners. The instructions and object placements vary across episodes. Cluttered: objects are placed close to each other in front and behind the robot, leading to OOD spatial layouts. The instructions and object placements vary across episodes.
We conduct experiments on the baselines and the proposed method, and measure their success rates in the real world. The results are summarized in Table I.
We run the whole training pipeline with three different random seeds to train low-level, the teacher and the student, and measure the performance of the final student policy in real, in terms of accuracy and efficiency:
  1. 1.
    Cumulative Subtask Success Rate: the success rate up to the specified subtask from the beginning of the episode.
  2. 2.
    Episode Time: the time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$s. For failed episodes, we use $t_{\rm max}$ as episode time.
Finally, we report the mean and standard deviation across the seeds as the final aggregated metric value in Table I.
Table I shows how important each of the interventions used contribute to the final successes. No Arm Retract baseline performs well for the initial two subtasks of Search and MoveTo, but not for the later subtasks. Even though half of the times the robot can move to the basket carrying the cube (a cumulative success rate of $28.3\%$), it only successfully drops the object into the basket less than one fifth of the cases (a cumulative success rate of $5.0\%$). The main reason is that without arm retract reward, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm. This leads to more arm shake, cube slipping out of the fingers, motor failures, and makes it harder for the arm to raise up to search for the basket and to hover above the basket to complete DropInto. No Perturb baseline has the lowest grasping success rate. Without perturbations (object, arm mount and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot. This is likely because during simulation training, the policy can get very efficient at approaching the target object and stopping just at the success boundary of the MoveTo subtask, so that the Grasp subtask usually starts with very similar locations of the target object relative to the robot, while in the real world such clear cut positioning is rarely seen. No Visual Aug’s performance on the initial Search subtask is lower than the that of the above baselines. The main reason is that without visual augmentations, it can be distracted by background objects irrelevant to the task, i.e. false positive detection of the target. This shows that visual augmentations are important for improving the robustness of the robot’s vision module. Distillation Only baseline simply imitates from the highly successful teacher policy, and can indeed achieve a relatively high success rate. However, its performance is not consistent across random seeds. Our full method (SLIM) uses distillation-guided RL, and results in consistently high success rates across seeds. It shows that distillation loss alone is not enough for robust policy learning. Additionally, task completion is also slower than the full method which additionally uses RL loss to optimize the policy.
Our full method SLIM achieves the highest subtask and full task success rate among all the methods, and is also the most efficient, taking only 43.8 seconds to complete the full task on average.
[…]
Figure 9: Real-world Performance of SLIM across Scene Variations. Left: images showing the differences for evaluation. Right: the full task success rate across difference scenes.
VIII-E Generalization to Different Real-world Scenes
In addition to the standard scene used in Table I, we further run real-world evaluations of SLIM under more scene variations (c.f. Figure 9 left): Outdoor, Carpet-Room, Kitchen and Lobby. For the first three scenes, we use the standard spatial layout (c.f. Figure 8). For Lobby, we use the Cluttered spatial layout, which is different from the training (Standard) layout. Results as summarized in Figure 9 are close to the Standard layout or a bit lower due to being out of distribution. This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, background, and floor materials and textures.
Furthermore, our qualitative evaluation showed that the robot also generalizes well to scenarios where:
  1. 1.
    task objects are randomly scattered on the ground (e.g., last two rows in Figure 1),
  2. 2.
    distractor objects are present (e.g., last row in Figure 1),
  3. 3.
    human interrupts the task progress, or
  4. 4.
    objects to grasp have slightly different shapes.
We refer the reader to our demo videos for more emergent generalization behaviors. In general, the robot is quite robust to the real-world scene complexity.
VIII-F Comparison with Simulation Performance
Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variability in the RL training of the teacher policies. Additionally, because of the sim-to-real gap, real deployment can differ from simulation as well.
Figure 10 shows simulation performance with the real performance for our methods, so we know more clearly how each method helps the training process.
The Sim success rate for each method denotes the average full task success rate over 200 episodes run in simulation, averaged across the three random seeds. The simulation environment follows the same randomization setting as training, with the policy rollout being fully greedy. The Real success rate denotes the full task success rate of the policy in the real-world, following the Standard spatial layout as shown in Figure 8.
Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasp, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up, creating an exploration bottleneck for RL training. Real world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: The tip of the arm tends to shake more due to motor backlash, hard to model in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overload. For No Perturb, while one might expect perturbations to complicate RL learning, their inclusion results in a much higher performance even in simulation. Without perturbations, the policy seems to converge to suboptimal solutions maybe due to a lack of exposure to a sufficiently large range of setups. Without visual augmentations (No Visual Aug), simulation task is easier to solve, thus simulation success is artificially high, but real world performance suffers when compared to the full method. RL training loss in the full SLIM method increases success in simulation (over the imitation only Distillation Only) and achieves the highest success rate in real. It’s interesting to note that the RL loss seems to help the No Visual Aug baseline to achieve slightly better real world success than Distillation Only.
[…]
Figure 10: Full Task Success Rates of Different Methods (averaged across three seeds) in Simulation and Real.
[…]
Figure 11: Policy Trajectory Visualization. For visualization purpose, we removed background objects. We also visualize the ee trajectories. Trajectories corresponding to different subtasks are rendered with different colors. The ego-centric camera view is shown on the top-right of each image. The red colored arrows denote the direction of movement.
VIII-G Policy Behavior Visualization
The behavior of the policy is shown in Figure 11. To make the visualization clear we did not add background objects in the simulation for this episode. The robot start with a randomized position and yaw within a region (Start phase). A red cube and a green cube as well as a red basket and and blue basket are used. Their positions are randomized around the Standard layout and their yaw angles are also randomized. Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube). Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f. Figure 11-Search). Once the target object is in view, the robot will move towards it (Figure 11-MoveTo). Once the robot moves close to the target object, it will start to grasp (Figure 11-Grasp). During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it. After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper. As shown in Figure 11-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory). At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE trajectory in Figure 11-SearchWithObj. Once the target basket is in view, the robot will move towards it (Figure 11-MoveToWithObj) and drop the cube in gripper into that basket (Figure 11-DropInto).
VIII-H Failure Modes
We also recorded common failure cases for SLIM. In one case, when grasping the quadruped stopped too late and kicked the cube far away. In the other case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach. DropInto sometimes fails because of either early dropping right outside of the basket, or the quadruped kicking the basket away right before dropping when backing towards the basket. In rare cases where the gripper holds only the upper half of a cube, the robot tends to get stuck at hovering over the basket while holding the cube, because the cube is completely blocked by the gripper fingers in the wrist camera view. See our failure demo videos for more details.
IX Limitations
Although not the main focus of this work, the current form of low-level policy introduces several limitations. One limitation is the applicable type of terrains. Currently the low-level policy is only trained on flat terrain in simulation. Although it shows the potential to work across different types of grounds (e.g. concrete, tile floor, carpet, mulch etc.), it does not push the low-level policy learning to its full potential. We can potentially leverage the recent advances in locomotion policy training [62, 7] to further expand the ability and applicability of SLIM across more diverse types terrains, e.g. stairs, ramp etc.
Another low-level-related limitation is that the current form of low-level policy is not a wholebody policy. Although the current form of low-level policy is sufficient for the task considered in this work, by generalizing it to a wholebody policy [15], we can potentially further enlarge its potential workspaces.
As a pure sim-to-real approach that is trained from scratch, we did not leverage any pre-trained foundation models. As a result, we did not address large amount of visual concepts and language variations and instead focused on the right amount of variations that are both practical and representative enough.
There are also a few limitations on the robot system side. Currently for the visual perception, we only used the RGB image from the wrist-mounted camera, which is the minimum setting that is sufficient for the current work. To further improve the robot’s perception and therefore the overall ability in terms of navigation and obstacle avoidance, it might be necessary to incorporate additional sensors for perception such as depth, stereo and lidar etc. In additiona, this can also provide valuable information for the robot to improve its longer-term memories via SLAM-like mechanisms [57, 5].
X Conclusion, Discussion and Future Work
We present a complete legged mobile manipulation robot system for solving language-instructed long horizon task in this work. The policy is trained via the sim-to-real paradigm, together with a progressive policy expansion-based teacher policy for solving the long-horizon task, followed with a distillation-guided RL approach for student visuomotor policy learning. We further identified and reduced a number of crucial sim-to-real gaps. Experimental results compared with a number of baseline methods verified the effectiveness of the proposed SLIM method. Moreover, real-world testing of SLIM across different scenes and spatial layouts shows that the proposed method can generalize well, further demonstrating the effectiveness of the proposed whole pipeline.
For future work, we would like to address the limitations discussed in Section IX, including further expanding the applicability of SLIM across more diverse scenarios by improving locomotion policy training [62, 7, 15]. Extending the visual and language diversity is another future work. Equipping the robot with additional sensors and enable it with SLAM-like ability [57, 5] is also an exciting future direction.
Appendix A RL Formulation
A-A Subtask Boundaries, Rewards and Behavior Priors
Subtask Boundaries.
Since we decompose the full long task into a sequence of subtasks, in practice, we need to decide the task boundaries in order to properly divide the full task. In the sequential decompositional case as described in Eqn.(6), the task boundaries $\tau^{i}|\tau^{i+1}$ between two adjacent subtasks $\tau^{i}\rightarrow\tau^{i+1}$ can be defined by the success condition $\mathcal{S}(\tau^{i})$ of the subtask $\tau^{i}$ (i.e. $\tau^{i}|\tau^{i+1}\equiv\mathcal{S}(\tau^{i})$), since the success of a subtask $\tau^{i}$ naturally leads to the subsequent subtask $\tau^{i+1}$.
We provide the success conditions of each subtask (therefore the subtask boundaries) in the sequel:
  1. 1.
    $\mathcal{S}$(Search): succeeds when the target object is in the view of the wrist-mounted RGB camera;
  2. 2.
    $\mathcal{S}$(MoveTo): succeeds when the target object enters a workspace of the robot within arm’s reach;
  3. 3.
    $\mathcal{S}$(Grasp): succeeds when the target object is in gripper and being lifted above the ground higher than 5CM and lasts at least two timesteps;
  4. 4.
    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
  5. 5.
    $\mathcal{S}$(MoveToWObj): succeeds when the target object is in gripper and the target basket enters a workspace of the robot within arm’s reach;
  6. 6.
    $\mathcal{S}$(MoveGripperToWObj): succeeds when the target object is in gripper and the grippers moved to the top of the target basket;
  7. 7.
    $\mathcal{S}$(DropInto): succeeds when the target object enters the target basket.
Rewards.
We summarize the description of the subtask reward in Table II. The task reward represents the sparse task success ($+1$) or failure ($-1$) reward. For timeout, we used a reward of 00. Apart from the sparse task reward, we also incorporated some shaping rewards as shown in Table II. Most of the shaping rewards are delta-distance based shaping reward, encouraging two position of interest to getting closer to each other. Arm retract reward is also an instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f. Figure A.13). Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position. Alignment reward is also an instance is delta-distance based shaping reward, where the distance is defined based on the cosine between the left and right finger-tip-to-object vectors. Keep-grasping reward is a contact based reward issuing a penalty of $-0.1$ is the contact between gripper fingers and the object is lost.
TABLE II: Rewards and Behavior Prior for Subtasks
Subtask Reward Behavior Prior
Search task rotational
MoveTo task + distance
Grasp task + distance + alignment stationary
SearchWObj task + arm-retract rotational
MoveToWObj task + distance + arm-retract + keep-grasping
MoveGripperToWObj task + distance + keep-grasping stationary
DropInto task + distance stationary
Behavior Priors.
Another benefit of task decomposition is that we can easily incorporate different behavior priors to the policy. We used two types behavior priors in this work:
  • rotational: the forward velocity command ($\mathbf{c}[0]=0$) to the quadruped is kept as zero and therefore enable the quadruped to move in a rotational fashion. This is used as the behavior prior for search related subtasks (Search, SearchWObj).
  • stationary: both the forward and angular velocity of the command to the quadruped are hardcoded as zero ($\mathbf{c}=0$). This encourages the robot to be stationary. It is used in manipulation related tasks (Grasp, MoveGripperToWObj, DropInto).
In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g. the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior). The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution and being consistent with subtasks without these priors.
Hierarchical RL.
As typical in standard hierarchical RL stuctures, the high-level policy and low-level policies are running at different frequencies. We refer one high-level inference step as a re-planning step. The high-level re-planning period is $5$, meaning the high-level policy (teacher or student) run once every $5$ low-level steps, and the same high-level output is used across this $5$ low-level steps re-planning period for the low-level policy.
Following this structure and given a frozen low-level policy, we are interested in training the high-level (teacher/student) policies. Because of this, the RL is essentially running at the high-level. Therefore, the per-(low-level-)step reward during each re-planning period is accumulated as the step reward for one high-level step, i.e.
$r_{\rm high\_level\_step}=\sum_{i=1}^{5}r_{\rm low\_level\_step}^{i},$ (7)
and this accumulated reward $r_{\rm high\_level\_step}$ is used as the per-step reward signal for high-level RL training.
A-B Low-level Task and Reward
Low-level Tasks.
Details of low-level training tasks/commands and corresponding rewards are summarized in Table III.
TABLE III: Low-level Training Tasks and Rewards
Task/Command Name Sample Range Reward
forward velocity $(-0.5,1.2)$ m/s $e^{-\frac{|v_{\text{forward}}-v_{\text{target}}|}{\max(v_{\text{target}},0.5)}}$
angular velocity $(-1.0,1.0)$ rad/s $e^{-|v_{\text{yaw}}-v_{\text{target}}|}$
stay stationary with probability $0.1$ $\begin{cases}-1&\text{if any foot lifted}\\ 0&\text{otherwise}\end{cases}$
balance and survival
tilt thresh: $0.6$
height thresh: $0.26$
$-(\mathbbm{1}_{\text{over-tilt}}+\mathbbm{1}_{\text{falling}})$
$-0.3\cdot\mathbbm{1}_{\text{killed}}$
gait following (optional)
Low-Level Domain Randomization.
The domain randomization setting for low-level training is summarized in Table IV.
TABLE IV: Low-level Domain Randomization Parameters
Env Params Sample Range
base payload $(-0.5,3)$
torso mass center $(-0.15,0.15)$ for each dimension
feet friction $(0.25,1.75)$
feet softness (solimp[2]) [^3] Only the third entry of the solimp parameter of Mujoco is randomized. It is roughly proportional to the size of the deformation of its feet when the robot stands still on the ground. $(0.005,0.02)$
griper payload $(0,0.1)$
Kp of leg joints $(0.7,1.3)$
actuator gain of arm joints $(0.7,1.3)$
bumpiness of the ground $(0,0.15)$ for $90\%$, 00 for $10\%$
delay of sensors $(u-0.002,u+0.002),\ u\in(0.003,0.013)$
Appendix B Model Details
B-A Teacher High Level
The teacher network $\Pi\!=\!\{\pi^{k}\}_{k=1}^{K}$ is a set of sub-policies (referred to as policy set), as shown in Figure 3. The effective size of the policy set is $K\!=\!8$ (including the Idle subtask).
Each policy net $\pi^{k}$ (for subtask $k$) is implemented as as a 3 layer MLP, with the layer size as $512$ ($[512,512,512]$). A squashed Normal projection layer (a layer for generating Normal distributions) is used as the output layer for each policy net following the standard practice [19], generating the action distribution conditioned on the current input.
The critic net for each subtask uses the same MLP structure and with the size of ($[512,512,512]$). Frame stacking is used with a stacking size of $5$.
The re-planning period is $5$, meaning that the high-level policy predicts once every $5$ low-level steps.
B-B Student High Level
Maps Prediction U-Net.
An illustration of the U-Net [42] for predicting segmentation and depth maps is shown in FigureA.14. We will talk about the architecture details by input modality below.
  1. 1.
    Language. Due to the simplicity of our language instructions, we directly encode each instruction into a sequence of byte tokens. We set the max sequence length to be 100, padding shorter sequences with zeros while truncating longer sequences beyond this limit. Then each byte token is converted into an embedding vector of 128 dimensions by looking up a learnable embedding table. The sequence of language embeddings, with learnable positional encodings, is fed to an instruction encoder transformer [48] with self-attention. The transformer has 2 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 64 and a value dimensionality of 64. The MLP for generating the residual output has 128 hidden units. Finally, the output embeddings are average pooled to generate a compact instruction-level embedding of 128 dimensions
  2. 2.
    Proprioceptive state. We use a temporal stack of $N=5$. Different proprioceptive observations are simply flattened and concatenated together to produce a 1D robot state vector.
  3. 3.
    Vision. We use a temporal stack of $N=5$ RGB images for vision. The images are stacked along the channel dimension and fed to a U-Net encoder containing 6 convolutional blocks. Each block has a convolutional layer $(filters,kernel\_size,strides,padding)\triangleq(c,k,s,p)$, an instance norm layer [47], a ReLU activation, and a 2D max pooling layer (only starting from the second block). The convolutional layers are configured as $c=[32,64,64,128,128,128]$, and $k=3\times 3,s=1,p=1$ for all blocks. We use FiLM [39] conditioned on the instruction embedding to modulate each intermediate convolutional output, before it is fed to the next convolutional layer. After getting the output from the encoder, we concatenate it with the robot state vector and the instruction embedding, and feed it to the non-skip MLP. The MLP has two hidden layers of 512 units. Finally, the output from MLP is projected back to the output space of the encoder, and is fed to the U-Net decoder to produce the predicted map. The map has $M+1$ channels, where $M=4$ is the predefined maximal number of segmentation classes, and the remaining one channel is for depth prediction.
The maps prediction U-Net is trained by groundtruth segmentation and depth in simulation.
Policy Inputs Encoder.
An illustration of the policy input encoder network is shown in Figure A.15. The network is responsible to digest predicted segmentation&depth maps and proprioceptive observations, and produce a compact representational embedding for downstream RL networks. Again, we use a temporal stack of $N=5$ for both segmentation and depth map inputs. After converting each segmentation map to a one-hot representation, we stack all maps along the channel dimension, resulting in an input of $5\times(4+1)=25$ channels. The input is fed to a ResNet [21] of 5 residual blocks, each block with 64 filters, a kernel size of $3\times 3$, and a stride of 2. The output from the ResNet is spatially flattened, resulting in a sequence of vision tokens. On the other hand, we use an MLP with one hidden layer of 256 units to project the concatenated proprioceptive observation into a token embedding of 256 dimensions. This proprioceptive token is appended to the vision token sequence which is fed to a representation transformer with self-attention. The transformer has 5 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 128 and a value dimensionality of 128. The MLP for generating the residual output has 256 hidden units. Finally, we take the transformed output corresponding to a special learnable CLS token as the final output of the entire policy input encoder.
Critic and Actor Networks.
Both the critic and the actor have three hidden layers of $(512,512,512)$. The actor outputs a squashed Gaussian distribution as in SAC [19]. Note that the RL and distillation gradients will be propagated all the way back to the policy input encoder for learning the representation.
Stationary Bit.
The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table II). With this prior, the corresponding locomotion command distribution always has a zero mean and a small std. Taking argmax of the distribution, the teacher policy generates exactly zero locomotion command and thus the quadruped can remain stationary. However, if we directly distill this distribution into the student’s counterpart, it is almost impossible to guarantee a zero mode, due to the continuous nature of the distribution. Thus we employed a sationary bit trick to solve this issue. Specifically, we create an extra binary action distribution for the student for predicting whether the quadruped should be stationary or not, which is determined by checking if the locomotion command of the teacher is exactly zero. We then add a new prediction loss for learning this stationary bit, while masking out the KLD loss (only for the two locomotion command dimensions) if the stationary bit is true. For student inference, whenever the stationary bit distribution outputs 1, the locomotion command $\mathbf{c}$ will be overwritten to be zero.
B-C Low Level
For the low-level training, we used PPO [44] with a regularized online adaptation module Fu et al. [15].
Policy Input and Locomotion Command.
As shown in Figure 2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g.. The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning. During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table III. The arm pose is also adjusted periodically within each episode, to make the low-level policy’s command following robust against arm movement. For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm. Empirically, we discover that it might trigger some shakiness behavior in certain cases and therefore we only used $\mathbf{s}_{\textrm{leg}}$.
After training, the low-level policy can be used to follow commands issued by an external command issuer, e.g. either a higher level policy (as in SLIM and all the other baseline except Human Teleop), or human tele-operator (as in Human Teleop) in this work.
Actor and Value Networks.
Both the actor and value networks have three hidden layers of size $(256,256,256)$. The actor network outputs a Beta distribution instead of the typical Gaussian. For numerical stability, we set the minimum value of the shape parameters of the Beta distribution to $1$.
Online Adaptation Module for Sim-to-Real Transfer
. Similar as in [15, Figure 2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder. The privileged encoder takes privileged information as input and predicts a latent representation of environment extrinsics, while the adaptation encoder predicts the same latent representation depending only on robot sensory observations. Both encoders are trained to match each other and jointly with the RL objective. The privileged encoder network has two hidden layers of $(128,64)$ and the adaptation encoder has hidden layers of $(256,128)$.
B-D Human Teleoperation
To show that we’ve made reasonable efforts to make Human Teleop work, we list its failure modes. 1) The most common failure mode is task timeout. This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time to search and re-grasp. Failing to grasp fully can be due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision. 2) The arm can enter a twisted pose due to poor IK solutions, despite having already tuned IK as we can.
Human teleoperation is in general less accurate and slower than RL policy, due to Low level policy shaking, and human operator fatigue. Besides being slower, human teleoperation seems to be less energy efficient than our RL policies. When RL learned high-level policies can run smoothly on a particular low level policy, human teleoperation on the same low level policy seems to trigger over-current protection of Go1. This often happens when trying to raise an over-extended arm, causing Go1 to lose control of its leg motors. In evaluations, we ignore these Go1 hardware failures, and let the human operator retry for free.
B-E Model Parameter Count
We summarize in Table V the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
TABLE V: Number of Trainable Parameters.
Component Approx. Param. Count
Low Level $3$M
Teacher High Level (all subtasks) $10$M
Student High Level $16$M
Our deployed policy (“Low Level” + “Student High Level”) has a total parameter count of less than 19M.
Appendix C Important Training Hyperparameters
Below we list some important training hyperparameters of SLIM.
Teacher High Level.
We used four 3090 GPUs for training, with a batch size of $2000$, a learning rate of $1\times 10^{-4}$, gradient clipping of $0.1$ with the Adam optimizer, $120$ parallel environments, and training iterations of $4$M (corresponding to about $60$M environment steps). The total training wall-clock time was about 1 week.
Student High Level.
We used 60 parallel environments to collect rollout data for training the student, where the data was stored in to a replay buffer with a capacity of $600$k environment steps. The full policy was trained on four 3090 GPUs for about 2 weeks, for a total of $1.3$M iterations. This amounts to roughly $40$M environment steps in total. We used a batch size of $64$, a learning rate of $1\times 10^{-4}$, gradient clipping of $100$, and the Adam optimizer. For the mixed rollout schedule, we kept the probability $\beta$ of sampling from the student policy as $0.5$ until the $750$K-th iteration, and linearly increased $\beta$ so that it reaches $1.0$ at the $1.25$M-th iteration. The distillation loss weight $\alpha$ in Equation 5 was set to $0.01$. The fixed modal dispersion $\sigma$ for distilling the teacher policy was $0.05$.
Low level.
We used 5000 parallel environments to roll out and train the low-level policy. The policy was trained on one 3090 GPU for about three days for a total of 4B environment steps. We used a batch size of 5000. The Adam optimizer had an initial learning rate of $2\times 10^{-4}$ and a step decay to its $1/5$ after $70\%$ of the total training progress.
Appendix D More Sim2Real Techniques
D-A More Dynamics Gap Reduction Techniques
Minimizing Tracking Error via PID Control.
An arm joint tracking error $\mathbf{e}_{\textrm{arm}}$ is the positional difference between the arm command of the previous timestep and the joint positions of the next timestep,
$\mathbf{e}^{t}_{\textrm{arm}}=\mathbf{s}^{t}_{\textrm{arm}}-\mathbf{a}^{t-1}_{\textrm{arm}}.$
Intuitively, we know that if the tracking error $\mathbf{e}_{\textrm{arm}}$ is equivalent between sim and real for all encounterable states, then there is essentially no dynamics gap. With this, one solution would be to match the tracking errors as closely as possible through accurate embodiment modeling with low-level PD control tuning. In practice, achieving this alignment for the entire state distribution proves to be nontrivial given both the nonlinearity of manipulator dynamics in conjunction with numerous unmodeled effects such as wear-and-tear, motor backlash, etc.
Rather than matching non-zero tracking errors, another approach is to simply achieve near-zero tracking errors in both sim and real. In other words, whatever arm command is outputted, the resulting arm position should be roughly achieved $\mathbf{s}^{t}_{\textrm{arm}}\approx\mathbf{a}^{t-1}_{\textrm{arm}}$. In fact, prior studies have shown that tracking error minimization is one of the most reliable strategies for enabling successful sim2real transfer [3]. To do so, we use a PID joint position controller for our manipulator in both simulation and in real, with a small joint position change limit $z$ (Section III-A) to ensure minimal tracking error for a wide variety of pose configurations.
Though the leg joints also possess a dynamics mismatch, we note that policy’s task success is minimally impacted by sim-to-real locomotion differences. Therefore, we opt to simply use PD position control for the legs.
D-B More Vision Gap Reduction Techniques
No Camera Depth Data.
For our setting, language-guided visual recognition in the RGB space is the key to finish the task. Without pretrained vision models, this sim-to-real transfer in RGB has to be supported by our method. Given this, we choose not to use camera depth data because this will introduce an extra sim-to-real gap, which has to be reduced with denoising or hole-filling techniques [23]. We observe that depth can be reasonably estimated from stacked RGB images, even with monocular vision.
Color Modeling.
Vision sim-to-real, especially for colors, is notoriously difficult [41]. We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow. These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training. Color samples are shown in Figure A.12.
Figure A.12: Color samples. Each color had 1000 randomly sampled pixels used to estimate the distribution in simulation.
Appendix E Robot Implementation Details
E-A Control Latency Management
Figure A.13: Asynchronous System Architecture. Note that “Onboard*” refers to the low-level controller frequency being unknown due to the use of closed-source software. Furthermore, when using the WidowX’s PID controller, we also take advantage of Dynamixel’s onboard profiling for smooth control.
We measured the latency along the full control cycle, including motor sensor reading ($\sim$3ms), camera observation (30$\sim$40ms), network connection ($\sim$0.2ms with cable and $\sim$5ms on Wi-Fi), and model inference (20$\sim$40ms for the high level  [^4] Later with TensorRT, we were able to reduce this latency to around 10ms. However, the experiments in this paper were conducted without this option back then., and $\sim$5ms for the low level). We implement the following methods to address such delays as much as possible, while properly modeling them in our simulator.
  1. 1.
    Asynchronous low-high inference: most of the model latency comes from the high-level inference, which takes around 20 ms, but can sometimes peak to 40 ms. This could delay the low-level policy (50HZ) considerably and make the robot falter, if the two levels are executed in sequence. Alternatively, in the asynchronous mode, the low-level policy will reuse the previous locomotion command if the new one has not arrived yet, and thus still maintain its 50HZ frequency. This leads to a more stable gait of the quadruped.
  2. 2.
    Ethernet cable connection to the robot: Wi-Fi connection is more convenient than cable. However, we find the Wi-Fi connection between Go1 and the laptop to be unstable, sometimes spiking to a latency of 100ms, which is unacceptable for low-level motor control. It is possible to have a Wi-Fi solution with more powerful and stable hardware, or to run low-level inference on Go1 instead of over Wi-Fi, but we simply rely on cable connection in this work.
  3. 3.
    Sensor delay randomizations: motor sensor reading delay is sampled from $(3,13)$ms per episode with $\pm$2ms variation uniformly per control step. Camera observation delay is sampled from $(30,50)$ms uniformly per step with observations taken at 60HZ. Low-level control delay is sampled from $(5,7)$ms uniformly per step. About the high-level inference delay, we evaluated the policies trained without high-level inference delay and compared the performances with and without test time high-level inference delays. We did not find clear performance differences in terms of the success rate. Therefore, we did not use high-level control delay in policy training in this work. Meanwhile, it is an option that is provided in our implementation and can be turned on whenever necessary.
Figure A.14: The architecture of the student’s U-Net for predicting segmentation and depth maps.
Figure A.15: The architecture of the student’s policy input encoder that generates a compact observation embedding to be used by RL.
E-B Safety Measures for the Physical Robot
To reduce damage to the robot itself and the surroundings, we employ many strategies.
  1. 1.
    Bridging the sim-to-real gap is one important part of it, so we can have confidence after observing the robot’s behavior in simulation.
  2. 2.
    PID arm control, with reasonable action clipping, results in much smoother and more reliable arm motion, and reduces damage to the arm motors significantly.
  3. 3.
    The randomization of motor strength, payload, center of mass, arm mounting position, etc. allows the policy to adapt to the real world more easily, resulting in very accurate grasping execution, so the arm does not push into the ground or the target objects, which can lead to motor damage.
  4. 4.
    Control delay caused by policy inference time and communication delay leads to poor low level control. Go1 can topple over, and damage components mounted on top. For example, high level inference can sometimes take over 40ms on the laptop. We use asynchronous inference of high and low level to make sure high level inference does not prevent the robot to receive low level commands. We find that the latency of the Go1 Wi-Fi module can spike to 100ms at times, and decide to rely on cable connection instead. We also allow terminating deployment with a button press. When latency still spikes, e.g. when the deployment machine suddenly freezes or network cable disconnects, our driver onboard Go1 automatically stops the quadruped’s motion. This stopping is done using a scripted policy, and is not always safe, but it prevents damage to the hardware in most cases.
  5. 5.
    Additionally, to prevent the robot from wandering off and bumping into the surroundings, during training, we incorporate a rotational search behavior as explained in Section A-A during policy training, for all baseline methods and the proposed SLIM method.
Appendix F Real-World Experiment Details
F-A Containers and Graspable Objects
We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Figure A.16. The size of the basket is about $16{\rm cm}\!\times\!16{\rm cm}\!\times\!16{\rm cm}$. The size of the cube is about $2.5{\rm cm}\!\times\!2.5{\rm cm}\!\times\!2.5{\rm cm}$ and the weight of the cube is about $9{\rm g}$.
[…]
Figure A.16: Containers and Graspable Objects.
F-B Real-world Evaluation Protocols
We detail the protocols used for real-world quantitative evaluation in the main paper (Table I).
TABLE VI: Scene Setups for Real-World Quantitative Evaluations. Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$. Uppercase letters denote baskets and lower case letters denote the objects to be grasped. The color of an object is denoted both by the background color and the character (e.g. g denotes the green cube and R denotes the red basket.)
Scene setup Instruction: ”Drop x into Y” (x $\mapsto$ Y)
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
b $\mapsto$ R
b $\mapsto$ B
[…]
y $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
  1. 1.
    For each method, we run 20 episodes with different initial scene states;
  2. 2.
    We apply a time limit of 90 seconds for each evaluation episode.
  3. 3.
    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Figure 8, as it is easy to ensure accuracy of the scene setup in terms of spatial positions. In practice, we mark the positions for the robot at the center of the square. Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets. The size of the square is about $2\times 2$ meters.
  4. 4.
    We follow a scene initialization procedure as detailed in Table VI across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
F-C Visualization Predicted Segmentation and Depth Maps
[…]
Figure A.17: Maps prediction results for some key frames of a deployment session. Time flows from top to bottom, and each row corresponds to one time step. From left to right, the columns represent the RGB frame, the predicted depth map, the predicted segmentation mask, and our added note for better interpretation of different stages.
In Figure A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session. We emphasize again that the robot was deployed in this zero-shot scenario never modeled by our simulator, and no model finetuning was performed. Despite this, we can see that the prediction results are high-quality, given only a very short history ($0.5$ seconds) of a single wrist-mounted RGB camera and the robot proprioceptive state. While searching for the target cube, the perception model had to locate it with just a handful of pixels (first row). Right after the gripper holding and lifting the cube above the ground, the model correctly predicted that the target had switched from the cube to a basket (cube’s green color turned red; third row). Moreover, while searching for the target basket in the outdoor environment, the cluttered background did not confuse the depth and segmentation prediction (fourth row). The perception model was also able to correctly predict the segmentation mask for the robot itself (gripper fingers and quadruped in orange; fifth row). Overall, we find that in practice, this high-quality visual information bottleneck not only greatly improves the generalization ability of our RL policy, but also facilitates our debugging of the entire robotic system as it separates the errors of the perception model from those of the policy.