### v3

# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

Abstract
We present a low-cost legged mobile manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning purely in simulation. This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following task instructions, and a low-level quadruped locomotion policy, 2) a teacher and student training pipeline for the high level, which trains a teacher to tackle long-horizon tasks using privileged task decomposition and target object information, and further train a student for visual-mobile manipulation via RL guided by the teacher’s behavior, and 3) a suite of techniques for minimizing the sim-to-real gap.
   In contrast to previous works that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX-250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer. Trained fully in simulation, a single policy autonomously solves long-horizon tasks involving search, move to, grasp, transport, and drop into, achieving nearly 80% real world success. This performance is comparable to that of expert human teleoperation on the same tasks but significantly more efficient, operating at about 1.5$\times$ the speed. Finally, we perform extensive ablations on key techniques for efficient RL training and effective sim-to-real transfer, and demonstrate effective deployment across diverse indoor and outdoor scenes under various lighting conditions.
Figure 1: SLIM in Real. Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in terrain, background, distractors, and other environmental factors. These scenes are not available in simulation training. The subtask annotations at the bottom are added for understanding task progress and are not part of the input to the system.
I Introduction
Legged mobile manipulation combines a robotic manipulator with a legged mobile platform, enabling robots to perform a wide variety of complex tasks in diverse environments [16, 59, 41, 33, 43, 65]. Unlike stationary or wheeled manipulators, legged systems can adapt to uneven terrains, significantly expanding the scope of potential applications. Recent advancements in hardware and algorithms [16, 46, 39] have also made legged manipulators more accessible and cost-effective, encouraging their adoption in research and industry.
Despite its promise, legged mobile manipulation presents unique challenges. The expanded scope of the tasks, the long task horizon, the diverse scenes and terrains, and unstable legged base all compound the difficulty of achieving reliable performance.
One popular approach to tackle mobile manipulation is Imitation Learning (IL) from expert demonstrations [5, 18, 54, 4]. It is straightforward, leveraging human expertise to produce reasonable behaviors quickly with supervised training. However, IL relies heavily on large and comprehensive datasets to ensure generalization and robustness, making data collection resource-intensive, both in terms of human labor and hardware wear-and-tear. Moreover, its performance is upper-bounded by the quality of human demonstrations. These limitations become even more pronounced when addressing the complexity of legged manipulation.
Another paradigm in robotics learning is sim-to-real Reinforcement Learning (RL). This method bypasses the need for extensive real-world demonstration data by leveraging virtually infinite data available from simulation, ensuring robust policy training through diverse and comprehensive data [50, 26, 28, 46, 33, 11]. Furthermore, simulation allows safe exploration without physical wear-and-tear and is not constrained by the performance of demonstrations, unlike IL. However, the success of this paradigm hinges on tackling two key challenges: (1) effectively training RL policies in simulation over long horizons, and (2) addressing sim-to-real gaps, including those both in dynamics and vision. These challenges are further exacerbated by the additional complexities introduced by legged manipulators.
Based on sim-to-real RL, we present SLIM (Sim-to-Real Legged Instructive Manipulation), a standalone pipeline for training robotic policies entirely in simulation and deploying them zero-shot in the real world. SLIM employs a hierarchical policy structure that decouples task-dependent high-level visuomotor commands from low-level control (Section III-B). In high-level policy training, we utilize a teacher-student learning framework to enhance training efficiency, where a teacher policy leverages privileged information to guide the student. We train the teacher policy to solve long-horizon tasks via task decomposition and progressive learning (Section IV). We then train a student policy, conditioned on language and visual inputs, in RL guided by the teacher’s behavior (Section V). Together with a suite of carefully designed sim-to-real techniques (Section VII), including visual and dynamics randomization and low-level controller tuning, SLIM produces visuomotor policies that can maintain sim performance when transferred to real.
In this work, we implement SLIM on a Unitree Go1 quadruped with a top-mounted WidowX-250S manipulator and a wrist-mounted Intel RealSense D435 camera (only using the RGB stream), resulting in a large 19 degree-of-freedom (DOF) system. We focus on the long-horizon task consisting of multi-stage search, move to, grasp, transport, and drop into, achieving nearly 80% real-world success (Section VIII).
Our key contributions are as follows:
  1. 1.
    We develop a low-cost legged manipulation system SLIM. To the best of our knowledge, SLIM is the first end-to-end robotic system for solving long-horizon real-world legged manipulation tasks from sim-to-real RL alone.
  2. 2.
    We conduct an extensive set of real-world experiments with 400 episodes achieving human expert-level success ($\sim$80%) with $\sim$1.5$\times$ the speed of human expert teleoperation. We also demonstrate robust task success across a wide range of indoor and outdoor scenarios.
  3. 3.
    We identify crucial techniques for achieving successful sim-to-real transfer with inexpensive hardware, validated through extensive real-world ablations.
We plan to open source the code to facilitate future efforts on related research.
II Related Work
II-A Learning-based Quadruped Locomotion
Traditionally, quadruped locomotion has been tackled through classical control methods designed to follow hand-tuned gaits on flat ground [19, 9], dynamic rigid platforms [27], discrete terrain [2], and rough terrain [13]. Though impressive, such methods require significant human engineering efforts and can be brittle to environmental changes.
More recently, there has been explosive progress in using learning-based approaches for achieving quadruped locomotion. In particular, sim-to-real reinforcement learning (RL) has arisen as a robust solution showcasing impressive feats such as parkour [67, 8, 6] and high running speeds [38]. Other works have also focused on exploiting the strengths of legged locomotion over their wheeled counterparts through utilizing the strong inherent coupling of proprioception with egocentric vision [17, 1] and active estimation of the environment [40, 31, 34]. Furthermore, necessary tuning can be minimal compared to classical approaches [15], for instance, massively parallelized simulation has been shown capable of learning gaits in just minutes [46]. Finally, data-driven legged locomotion has shown great promise in its ability to zero-shot generalize to new morphologies [14], surfaces [39, 31, 34] and agile skills [56, 30].
II-B Legged Mobile Manipulation
Significant advancements in robust locomotion have further enabled researchers to push the boundaries of legged mobile manipulation. Building upon data-driven approaches, quadrupeds have been demonstrated pushing objects with their body [29], dribbling around balls [30], and manipulating objects using egocentric [55] or calve mounted grippers [32]. Still, the most popular setup has been the traditional top-mounted 6DOF manipulator design [16, 20, 33, 65, 59, 37, 41, 48]. In addition to significant workplace expansion, manipulators offer other nuanced benefits, such as assisting with balance [24] and serving as an intuitive interface for collecting human demonstrations [20]. Classical approaches to solving top-mounted legged mobile manipulation consist of designing wholebody controllers [48] or even combining locomotion learning with model-based manipulation control [37]. Others simplify the need for low-level controllers by using high-level API calls provided by their respective quadruped robots to achieve graspability-aware policies [65], navigational pick-and-place [59], and language-conditioned mobile manipulation [43]. Complex tasks have also been achieved by leveraging teacher-student training setups in simulation for grasping [33] and door opening [66]. Finally, to expand the robot’s workspace even further, wholebody loco-manipulation has arisen where leg joints are actuated in a way to assist manipulation [16, 33, 41].
Most closely related to our work, Liu et al. [33] introduces a visual whole-body control approach (VBC), which trains a sim-to-real visuomotor whole-body loco-manipulation policy using a high-low hierarchical model and teacher-student training setup for the object pickup task, similar to our design. The system can grasp an impressive set of diverse objects by leveraging whole-body manipulation with capable hardware. VBC takes a few shortcuts to simplify the task. VBC requires the user to manually click on the target object to perform initial segmentation using a third-party vision model. This segmentation is then used to track the object through a third-party tracking model, which requires the object to stay within the camera view the whole time. Additionally, VBC uses a scripted policy to put the object into the basket on the back of the quadruped, once the object is lifted, making the task scope more limited and the task horizon much shorter. In comparison, SLIM offers the follow advantages:
  1. 1.
    SLIM is a complete and self-contained system. All modules–vision, high-level visual policy, low-level motor control–are trained in our simulation environment integrated within a single framework without any third party modules, resulting in a lean 19 million parameter neural network system, allowing both minimal latency and full autonomy.
  2. 2.
    SLIM is fluid and intuitive to use. Given a language instruction, SLIM is able to rapidly accomplish tasks in a wide variety of scenes.
  3. 3.
    Finally, SLIM is practical. SLIM can accomplish complex long-horizon real-world tasks autonomously from the beginning to end, as apposed to some of the existing methods that requires human involvement (e.g. manually clicking for segmentation) and can only automate part of the task with restrictions on the feasible starting state distribution [33]
III Architecture Overview
[…]
Figure 2: Hierarchical Framework and Visuomotor Policy Pipeline (Left). Given language instruction and sensor inputs, the high-level policy generates a set of two control signals: 1) the arm and gripper control signals, and 2) the locomotion command. The arm control signals are directly passed to the arm driver, and the locomotion command is passed to the low-level policy to control the leg joints of the quadruped. Low-Level Training (Top Right), and High-Level Teacher and Student Training (Middle and Bottom Right). SLIM training is divided into three sequential stages. First, the low-level locomotion policy is trained via RL to follow a sampled linear and angular velocity command. Second, the high-level teacher policy is trained via RL with privileged low-dimensional state input to solve the long-horizon task. Finally, the student visuomotor policy is trained by distilling the teacher behavior while maximizing task rewards, using visual, sensory, and language instruction as input. Both the teacher and student policies command the same frozen low-level policy produced by the first stage. The teacher is only run in simulation and the student can be deployed in real.
Fig. 2 illustrates the hierarchical (high-low policy) framework (left) and the teacher-student training pipeline (right) used in SLIM. Given language instruction and sensor inputs, the high-level policy generates intermediate control signals that will be used as input to the low-level policy. A low-level locomotion policy is trained to track a specified locomotion command via RL and is frozen after training. Then a high-level teacher policy is trained with privileged information via RL to solve the long-horizon task, and a student visuomotor policy is trained by distilling the teacher policy while maximizing task rewards, conditioned on visual inputs and language.
III-A Observations and Actions
Observations.
The robot observation at each time step consists of three components:
  • $\mathbf{l}$
    A tokenized instruction vector of length $L$.
  • $\mathbf{o}$
    A temporal stack of RGB images from a single wrist-mounted camera with a shape of $(N,H,W,3)$, where $N$ is the stack size, and $H$ and $W$ are the image height and width, respectively. The environment is always partially observed by the robot, and most of the workspace will not be contained in this camera view during the task.
  • $\mathbf{s}_{\textrm{r}}$
    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive readings of all joints. $\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive state (leg joint positions) and $s_{\textrm{arm}}$ the arm proprioceptive state (arm joint positions).
Privileged Observations (only available in simulation).
To achieve better learning efficiency, we will first train another policy using privileged observations, and then distill it to our robot policy conditioned on the standard observations. This guidance policy has a subtask id input $k$ in place of $\mathbf{l}$ and a lower-dimensional object observation $\mathbf{s}_{\textrm{p}}$ in place of $\mathbf{o}$:
  • $k$
    An integer in $[1,K]$ indicating which subtask the teacher is currently solving. $K$ is the maximal number of subtasks forming a long-horizon task.
  • $\mathbf{s}_{\textrm{p}}$
    A group of temporal stacks of privileged object features (e.g., positions, scales, categories, etc.), where each stack has a shape of $(N,Q_{m})$ and $0\leq m<M$ denotes the object index. $M$ is the maximal number of task related objects in the scene. Note that object features are also partially observed and their visibility is always determined by the robot’s camera field of view. The features are set to zero if they are not in the view. This is to ensure that the object visibility is consistent between the guidance policy and the robot policy, and avoid any information gap for policy distillation.
The privileged information can be extracted or computed from the underlying simulator state.
Actions.
The robot policy outputs a triplet of actions:
  • $\mathbf{a}_{\textrm{arm}}$
    A vector in $[-z,z]^{I}$, indicating the delta changes to the arm joint positions. $z$ is the maximally allowed joint position change for a control interval. $I$ is the number of actuated arm joints, excluding the gripper joint.
  • $g$
    A target gripper position in $[0,1]$.
  • $\mathbf{a}_{\textrm{leg}}$
    The target quadruped joint positions in $\mathbb{R}^{J}$. $J$ is the number of actuated leg joints.
In total, a complete action $\mathbf{a}$ has $(I+1+J)$ dimensions.
III-B Hierarchical Policy Structure
The overall system takes visual and instructional input and controls all joints of the quadruped and the arm, which involves a non-trivial training task with high-dimensional input and output. We use a two-level hierarchical policy to divide the complexity of visual mobile manipulation training from legged locomotion training, illustrated in the left part of Fig. 2, similar to [33].
The high level policy takes visual, instructional, and proprioceptive inputs, and outputs a locomotion speed command and arm control command. The low level policy tracks the locomotion command from the high level, using proprioceptive inputs to command the quadruped’s leg joints. The assumption is that given the locomotion command, quadruped locomotion control is largely independent of the high-level task semantics.
Intermediate Action.
The following intermediate action separates the high and low-level policies,
  • $\mathbf{c}$
    A vector in $\mathbb{R}^{2}$, containing the target forward and angular velocities for the quadruped.
With this, a high-level policy is defined as
$(\mathbf{l},\mathbf{o},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a}_{\text{hi}}\triangleq(\underbrace{[\mathbf{a_{\textrm{arm}}},g]}_{\textrm{arm and gripper control}},\underbrace{{\mathbf{{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}[}c{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}]}}}}_{\textrm{locomotion command}})\in\mathbb{R}^{I+3},$
while the low-level policy is defined as
$(\mathbf{c},\mathbf{s_{\textrm{leg}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{J}.$
Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$ and thus, does not need to process high-dimensional image and language inputs.
Low Level.
The low-level policy is a quadruped controller that generates joint position targets for PD control to follow a task-agnostic 2D linear and angular-velocity command $\mathbf{c}$ [39] and is trained in simulation using PPO [47]. An illustration of the training can be found in the top row of the right column of Fig. 2. It is worth noting that during low-level training, the same embodiment as shown in Fig. 2 (arm mounted on top of the quadruped) is used in low-level training. We randomly sample both command $\mathbf{c}$ and arm joint actions $\mathbf{a_{\textrm{arm}}}$ during training. Since the arm mounted on the quadruped can have varying poses during the task, this enforces the low-level command following ability to generalize to various arm configurations. That is, the quadruped base has to learn to keep balance while achieving locomotion commands regardless of the arm’s current joint positions and movements. We intentionally remove arm joint state and joint command from low level input, see details in Appendix B-C. Similar to most sim-to-real approaches, we also randomize the simulation environment during training. Besides widely-used domain randomization parameters as in [16], we further randomize the simulated delay of each sensor and foot softness of the robot to better adapt to variations in real deployment. Full details of the training setup and reward design can be found in Appendix A-C.
High Level.
The high-level policy is responsible for outputting the locomotion command $\mathbf{c}$ and the arm controls $(\mathbf{a_{\textrm{arm}}},g)$, given language instruction $\mathbf{l}$, proprioceptive state $\mathbf{s}_{\textrm{r}}$, and the current stack of RGB observations $\mathbf{o}$. As shown in Fig. 2, the command $\mathbf{c}$ is forwarded to the low-level policy, which then follows the received command for a number of time steps, before the high-level policy outputs the next command. For the manipulator, we opt to operate in joint space, as opposed to task space or end-effector pose, so we do not have to worry about inverse kinematics computation. Once the low-level policy is trained, it is frozen and used as a base controller by the high-level (teacher or student) policies. In the remainder of main text, we only talk about the high-level policies unless otherwise stated.
III-C Teacher-Student For High-Level Training
With the low level policy taking over the responsibilities of legged locomotion control, high level can focus on task dependent decision making. However, the high level policy needs to process high dimensional visual and language input, and complete the long-horizon task. This presents an enormous space for the policy to explore, and a challenge for efficient RL training. We adopt a teacher-student learning framework  [12, 33, 66] for efficient high-level training. The overall structure of the framework is shown in the middle and bottom right sections of Fig. 2. As shown, the teacher is trained purely with RL from privileged, structured, and low-dimensional inputs $(k,\mathbf{s_{\textrm{p}}},\mathbf{s_{\textrm{r}}})$ some of which can be obtained from the simulator but not easily in the real world. When a teacher policy is successfully learned, it is frozen and used to guide the student via behavior distillation. As the student will eventually be deployed in the real world, its inputs no longer contain privileged information. Besides policy distillation, we also use an RL objective to allow the student’s behavior to be shaped by the same set of task rewards used by the teacher, for a potentially more effective policy than using distillation alone (details in Section VIII-B, Distillation Only baseline). In the two sections below, we explain the teacher and the student policies in more detail.
IV The Teacher: Long-Horizon RL with Task Decomposition and Policy Expansion
The Teacher works with privileged object state information, thus, offloading the complexities of visual representation learning to the student, and only needs to focus on the following challenges of long-horizon task learning,
  1. 1.
    Continual exploration: For long-horizon tasks, there may be a number of intermediate milestones that must be sequentially achieved. Therefore, even after reaching an intermediate milestone, the teacher must continue exploring new frontiers in order to ultimately solve the entire task. Without carefully encouraging continual exploration, the agent can stop exploration early and settle on a suboptimal solution.
  2. 2.
    Loss of capacity and catastrophic forgetting: As the teacher progresses in training, it needs to cope with both the loss of capacity issue [35, 36], which hinders the network from continual learning, and the catastrophic forgetting issue, which could destroy skills that have already been acquired during the learning of new tasks.
Addressing these challenges are key for successful teacher policy learning. We propose to use task decomposition and further integrate with policy expansion (PEX) [64] to address these two challenges.
We first decompose the long horizon task $\mathcal{T}$ into $K$ subtasks $\{\tau^{k}\}_{k=1}^{K}$ with shorter-horizons. To incorporate the task decompositional structure in learning, we leverage the fact that privileged information is accessible to the Teacher and thus include the subtask index $k$ as part of the privileged observation. Subtask successes keep pushing the agent to learn new subtasks one after another, addressing the exploration issue.
Secondly, we solve the loss of plasticity and forgetting in long horizon task learning via a progressive application of the PEX scheme [64]. Originally designed for two-stage offline-to-online training, PEX uses dedicated networks for learning within each stage. This way, the trainer creates new network instances for continual exploration and learning of new skills without affecting any skills acquired in the previous stage. Here we apply this idea progressively to the decomposed long-horizon task with multiple subtasks, where the stages corresponds to the subtasks. We dub this strategy as Progressive PEX.
Because of policy expansion at each stage transition in Progressive PEX, it addresses the two challenges brought by the long-horizon task learning naturally: 1) it can achieve multi-stage continual exploration and the full ability of exploration is always ensured when entering a new stage; 2) since the policy for solving one subtask is encapsuled in a dedicated network, it mitigates the catastrophic forgetting issue for the already learned subtask policies. For a newly encountered subtask, the dedicated network that will be newly allocated addresses the issue of lost plasticity.
[…]
Figure 3: Teacher Policy Network Structure. The teacher network is a set of identical submodules $\{\pi^{k}\}_{k=1}^{K}$ gated by the subtask id $k$ (right side of figure). Each submodule $\pi^{k}$ is only trained with the experience when the agent is working on subtask $k$, taking privileged and proprioceptive input, flattening and concatenating them before passing it through an MLP (left side of figure). On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP. On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$. Given an input id $k$, the $k$-th component network is activated in the computation, i.e., $\Pi[k]\equiv\pi^{k}$.
We use $\pi_{\rm tea}^{k}(\mathbf{a}_{\text{hi}}|\mathbf{s})$ to denote the teacher sub-network that is responsible for solving its corresponding subtask $\tau^{k}$. When there is no confusion, we omit the subscript of the teacher network $\pi_{\rm tea}^{k}$ as $\pi^{k}$. The left side of Fig. 3 provides a graphical illustration of the teacher sub-network structures. Each sub-network takes low-dimensional proprioceptive and privileged observation as input and is responsible for learning to solve the corresponding subtask.
Progressive PEX works as follows. We initiate the exploration and learning with a single policy network [^1] We use the term policy network to represent all the networks that are required for learning a policy. For example, in the context of Actor-Critic formulation, it encompasses both the actor and critic networks. $\Pi=\{\pi^{1}\}$ that is responsible for learning to solve the initial subtask. Whenever a new subtask is encountered, a policy expansion operation is conducted, adding a new policy into the policy set, i.e., $\Pi=\{\pi^{1},\pi^{2}\}$. By doing this progressively for the all the $K$ subtasks, we get
$\Pi$$\triangleq\{\pi^{k}\}_{k=1}^{K},$ (1)
as illustrated in the right side of Fig. 3.
Based on this implementation, the training of $\Pi$ is closely related to multi-task RL [62, 49]. For each training iteration, we get a batch of samples containing the subtask indices from the replay buffer and use each sample in the batch for training the sub-network associated with the subtask index of that sample. In this work, we train the teacher policy using a multi-task variant of SAC algorithm [21, 62, 49]. The performance comparison between the method using a standard network structure and the Progressive PEX approach on solving long-horizon tasks is provided in Appendix J-B, which clearly shows Progressive PEX has a much stronger ability in long-horizon task learning.
V The Student: Policy Distillation Guided RL
The student needs to perceive the surrounding environment using RGB, motor sensors and language instruction. Because the student needs to be deployed in the real world, it cannot use any privileged information e.g. which subtask it needs to solve. Therefore, the student has to learn a single policy over the entire long-horizon task.
To tackle these challenges, we train the student by distilling the teacher’s behavior into a single task policy $\pi_{\rm stu}$. However, distillation on its own is likely to produce a policy that works reasonably well but it could lead to state drifting / out-of-distribution (OOD) issues when the policy is rolled out over a long horizon. There is also always an observation gap between the teacher and student, which may create difficulty in minimizing the distillation loss. Moreover, due to differences in model architectures and inductive biases, the student may need some adaptation of the teacher’s skills. Given these considerations, we choose to boost the student’s policy with RL under the same set of task rewards used to train the teacher (Fig. 4).
We train the student by modifying SAC [21] to incorporate the distillation loss properly. First, we use a mixed rollout strategy to generate replay data. When rolling out a new episode, with a probability of $\beta$, we will sample actions from the student policy, otherwise from the teacher policy. On the one hand, we wish to use the teacher to efficiently generate high-performing transitions along the long horizon. On the other hand, we want to keep exploring with the student’s own policy. Second, following SACLite [61], we remove the entropy reward from policy evaluation. Then for policy improvement, we replace the entropy term with the distillation loss and assign a fixed weight $\alpha$ to it:
$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{s}_{\textrm{stu}},\mathbf{s}_{\textrm{p}},k)\sim\mathcal{D}_{\text{replay}}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{s}_{\textrm{stu}})}Q\big{(}\mathbf{s}_{\textrm{stu}},\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha{\rm KL}\!\left(\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{\textrm{r}},\mathbf{s}_{\textrm{p}})\,||\,\pi_{\rm stu}(\cdot|\mathbf{s}_{\textrm{stu}})\right)\Big{]},\\ \end{array}$ (2)
where $\mathbf{s}_{\textrm{stu}}=[\mathbf{o},\mathbf{s}_{\textrm{r}},\mathbf{l}]$. To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std. for Gaussian). This results in a distillation loss that elegantly achieves two goals: imitation learning and entropy regularization.
[…]
Figure 4: Student Architecture Overview. Red dashed lines represent data flows that only exist in simulation training. Solid lines denote the inference process of the student policy.
When designing the student’s representation model, we specifically take visual sim-to-real gap reduction into consideration, because once the student is trained, it will be directly deployed in real without any finetuning. While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck (Fig. 4) to reduce the visual sim-to-real gap while achieving better interpretability.
The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space. This can leads to potentially better generalization. While there might be many choices for the vision bottleneck, from discrete latent codes [22] to canonical images [28], we choose to use a pair of segmentation and depth maps. There are mainly three reasons for this choice. First, we believe that they contain the minimal visual information required to achieve our task without incurring much task-relevant information loss. Second, ground-truth segmentation and depth maps can be efficiently obtained from the simulator for high-quality supervised learning of such a vision bottleneck. Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs. Note that only the robot and task-related objects appearing in the instruction will be predicted in the segmentation mask. All other irrelevant object pixels will be classified as “background”. After this, the segmentation and depth maps will be fed to the downstream policy for generating actions. Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [33, 59]. Finally, this design leads to interpretable visual representations, which is helpful for us to diagnose the visual sim-to-real gap.
VI Task Implementation In Simulation
Figure 5: Task Decomposition as Privileged Information (for teacher only). Given a task, we decompose it into a sequential series of primitive subtasks. As shown above, the task of finding an object A, grasping it, and then dropping into an object B can be decomposed into a sequence of subtasks. In simulation, the subtask boundaries and transitions are well-defined, serving as a form of privileged information to the teacher. The teacher can leverage this information, and use a separate policy to learn each subtask. This expertise is then distilled into the student, which is conditioned on both visual camera feed and language commands.
As mentioned previously in Section IV, SLIM employs a task decompositional approach paired with progressive PEX for training a teacher policy to solve the long-horizon task. In this work, we use a multi-stage search and pick-and-place task as our exemplar long-horizon task (Fig. A.16). This task will also serve as a concrete example for illustrating the task decomposition. Given a language instruction (e.g.,Drop the blue cube into the green basket.”), the robot needs to locate an object A (the blue cube), pick it up, and then drop it into a container B (the green basket). This task can be naturally broken down into the following sequential steps, where object A and B are specified by the language command itself:
  1. 1.
    Search: The robot must first search for object A using the RGB images from the wrist-mounted ego-centric camera.
  2. 2.
    MoveTo: After locating object A, the robot should approach it until it is within the reach of the arm.
  3. 3.
    Grasp: The robot attempts to grasp and pick up object A.
  4. 4.
    SearchWithObj: After picking up object A, the robot should then search for container B. Within this stage and those after, there should be some coordination between the arm and locomotion movement to facilitate subtask success while avoiding dropping the grasped object.
  5. 5.
    MoveToWithObj: After locating container B, the robot should approach it until it is within the reach of the arm.
  6. 6.
    MoveGripperToWithObj: The robot then moves its gripper towards container B until reaching an appropriate position over the top of the container.
  7. 7.
    DropInto: Finally, the robot releases its gripper and drops object A into container B.
An illustration of the full task decomposition can be observed in Fig. 5. Note that in this figure, we append an auxiliary Idle subtask, which encourages the robot to adjust its arm back to its starting position after finishing DropInto. This serves as an optional subtask and is used purely to facilitate subsequent deployments. Herein after, we do not include Idle in evaluations or discussions as it is unrelated to the task.
Task decomposition offers many benefits. For one, we can implement reward functions for each subtask, aiding modularity and ease of implementation. For all subtasks, we incorporate a sparse subtask success reward. Additionally, for those requiring moving the robot (or gripper) to a target region (e.g., MoveTo, Grasp, etc.), we incorporate an additional dense delta distance-based shaping reward. For SearchWithObj and MoveToWithObj, we also incorporate an arm retract reward (another delta distance-based shaping reward), where the arm is encouraged to stay close to a predefined target pose, shown in Fig. 6. This serves to promote good vantage and discourage suboptimal behaviors after grasping such as looking at the ground when searching. Its impact will be investigated in Section VIII-B (No Arm Retract baseline).
Finally, task decomposition offers the extra advantage of allowing behavior priors to be flexibly integrated into the policy. Such behavior priors can significantly increase sim-to-real task success as we will show later in Section VIII-D. More details on the subtasks, rewards, and behavior priors are provided in Appendix A-A.
[…]
Category Technique Description Reference Importance
Rating (1-3)
[…]
a deterministic behavior that fails in the real world
[…]
mismatch in the absence of system identification
[…]
for better generalization in real
[…]
VII Sim-to-Real Gap Reduction Techniques
In this section, we discuss crucial techniques for addressing sim-to-real gaps and achieving successful real-world deployment.
The dynamics sim-to-real gap is the mismatch between the transition function of the simulation versus that of the real world. It can be due to the mismatch in the physical properties of objects and motors, or by things that are not properly simulated, such as friction and backlash. This difference in dynamics can cause policies trained purely in simulation to fail when deployed in the real world, especially if the task requires accurate motor control, such as grasping a small object.
The second source of sim-to-real gap for image-conditioned policies is the vision gap. This gap is caused by an RGB distribution mismatch between simulated and real-world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios.
For brevity, we list techniques we deemed most relevant in Table I. Comprehensive details can be found in Appendix E.
VIII Physical System, Experiments and Results
VIII-A Physical Robotic System
For our robotic system, we use a Unitree Go1 with a top-mounted WidowX-250S manipulator. An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount. The RGB stream of D435 is used as the robot’s sole vision feed (Fig. 6). Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach. The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction. Finally, the base of the WidowX is raised a few inches to allow for easy access to the onboard Raspberry PI’s ports. Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around \$12K USD. With the non-edu version, total cost is even lower ($\sim$6K USD) and reflects a bit more closely the cost of mass produced parts..
Figure 6: Qualitative Overview of the Robot System.
TABLE II: Dimension values for the observations and actions of our robot. For the meanings of these notations, we refer the reader to Section III.
$L$ $H$ $W$ $N$ $D$ $K$ $Q_{m}$ $M$ $z$ $I$ $J$
$100$ $90$ $160$ $5$ $66$ $7$ $19$ $6$ $0.05$ $6$ $12$
[…]
Given our hardware configuration, the dimension values from Section III are summarized in Table II.
VIII-B Ablations and Baselines
We compare SLIM against the following baselines, where the first four are ablations and the last is a human teleoperator as a reference:
  • No Arm Retract: Remove the arm retract (raise) reward in RL training (Section VI).
  • No Perturbation: Remove arm control randomization, arm mount perturbation, and object perturbation (Section E-A).
  • No Visual Aug: Remove random background objects and spatial visual augmentations (Section E-B).
  • Distillation Only: Train student without the RL loss, keeping only the distillation and representation loss (Section V).
  • Human Teleop: An expert human teleoperator that provides locomotion commands and delta EE pose to the robot via a joystick. Delta EE is handled via IK. The teleoperator shares the same observation space as the policy, i.e., observing solely through the wrist camera stream, and is allowed to practice on the system for an hour. While not a baseline intrinsically, human teleoperation provides a valuable reference point. See Appendix I for more details.
We also tried another baseline No Distillation, which trains the student policy directly via RL without distillation from the teacher policy (i.e., RL for visuomotor policy learning from scratch). This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration. As such, it is excluded from the results.
VIII-C Metrics and Evaluation Protocol
To accurately evaluate policy performance, we run the entire training pipeline across three different random seeds for all methods, training from scratch the low-level, teacher, and student policies. Note that each method for a particular seed shares the same low-level policy. We report the following two key metrics when deploying the final student policy:
  1. 1.
    Cumulative Subtask Success Rate: The success rate up to each subtask from the beginning of the episode.
  2. 2.
    Episode Time: The time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$ s. For failed episodes, we use $t_{\rm max}$ as the episode time.
For the deployment environment, we use the Standard object spatial layout in an indoor Lobby scene, as shown in Fig. 7, to maintain repeatability across all methods and seeds. For task objects, we use cubes and baskets of different colors. For each method of each random seed, we roll out the robot’s policy for 20 episodes with varying object types and positions according to a consistent evaluation protocol, with 360 real-world episodes of experiments. More details on the physical objects and protocol can be found in Appendices G-A and G-B, respectively. In addition to the Standard spatial setup in Lobby, we also evaluate the robot in several other scene and layout combinations in subsequent experiments, with another 40 real-world episodes, leading to a total of 400 real-world episodes of experiments.
TABLE III: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
Method Autonomous Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
[…]
No Visual Aug 98.3±2.498.3.2.498.3$\pm$ 2.498.3±2.4 88.3±4.788.3.4.788.3$\pm$ 4.788.3±4.7 63.3±26.663.3.26.663.3$\pm$ 26.663.3±26.6 56.6±22.556.6.22.556.6$\pm$ 22.556.6±22.5 52.9±18.752.9.18.752.9$\pm$ 18.752.9±18.7
Distillation Only 93.3±4.793.3.4.793.3$\pm$ 4.793.3±4.7 71.1±33.271.1.33.271.1$\pm$ 33.271.1±33.2 60.0±38.960.0.38.960.0$\pm$ 38.960.0±38.9 50.0±35.550.0.35.550.0$\pm$ 35.550.0±35.5 61.2±25.561.2.25.561.2$\pm$ 25.561.2±25.5
[…]
VIII-D Results and Analysis
[…]
Figure 7: Scene Spatial Layouts in Lobby Environment. Standard: we put the robot in the center of a square space and objects including objects to be grasped and containers to drop the graspable objects into on the four corners. The instructions and object placements vary across episodes. Cluttered: objects are placed close to each other in front and behind the robot, leading to an OOD spatial layouts. The instructions and object placements vary across episodes.
All success rate and task completion results are summarized in Table III. As shown, we can see the negative impact of ablating away each technique on the final task success.
In particular, No Arm Retract performs fairly well up until Grasp, but then soon its performance drops off substantially. In fact, No Arm Retract possesses the lowest cumulative success rates past grasping across the board, with a sharp dropoff of approximately 30% when attempting to search the container. There are several reasons for this dropoff. For one, without the arm retract reward acting as a soft constraint on the arm’s position, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm. This leads to more arm shake, which in turn can result in the grasped cube slipping out of the fingers or even motor failures. Another key reason for search-based failure arises from possible low vantage search strategies, e.g., the camera is pointed toward the ground.
Intuitively, No Perturbation has the lowest grasping success rate. This makes sense as without perturbations of any kind (object, arm mount, and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot. This is likely because during simulation training, the policy can get very efficient at approaching the target object and stopping just at the success boundary of the MoveTo subtask, so that the Grasp subtask usually starts with very similar locations of the target object relative to the robot, while in the real world such clear cut positioning is rarely seen.
Performance of No Visual Aug on each of the Search subtasks suffer as the model sometimes gets distracted, having false positive detections of the target in the background. There’s a particularly large drop of roughly 25% when searching for the basket. This indicates that visual augmentations are important for improving the robustness of the robot’s vision module.
Finally, Distillation Only simply distills the highly successful teacher policy, and can indeed achieve a relatively high success rate. However, its performance is not consistent across random seeds. In comparison, SLIM uses distillation-guided RL, resulting in consistently high success rates across all seeds indicated by the low standard deviations. It shows that distillation loss alone is not enough for robust policy learning. Additionally, task completion is also slower as the RL loss allows the policy to further optimize. Overall, SLIM achieves the highest subtask and full task success rates among all the methods including Human Teleop, and is also the most efficient, taking on average 43.8 seconds to complete the full task, which is about 1.5$\times$ the speed of a teleoperation expert.
We provide some visualization of the policy behaviors and prediction outputs in Appendix H.
[…]
Figure 8: The same SLIM Policy across Scene Variations. Left: images showing the differences for evaluation. Right: the average full task success rate across 10 real-world deployment episodes for each of the scenes shown on the left.
VIII-E Generalization to Different Real-world Scenes
In addition to the standard scene used in Table III, we further run real-world evaluations of SLIM under more scene variations (left side of Fig. 8): Outdoor, Carpet Room, Kitchen, and Lobby-Cluttered. For the first three scenes, we use the same Standard spatial layout as before. For the Lobby-Cluttered scene, we use the Cluttered spatial layout (right side of Fig. 7), which is different from the Standard layout used during training. We conduct ten trials for each environment using the same SLIM policy.
All results for these four additional scenes are summarized in the right side fo Fig. 8. For the three novel scenes, we observe that success rates are all fairly close to the results from Table III ($\sim$78%). For the Lobby-Cluttered scene, we can see that the success rate is quite close to the Standard layout and is just a bit lower due to being out of distribution. This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, backgrounds, and floor materials and textures.
Furthermore, our qualitative evaluation showed that the robot also generalizes well to scenarios where:
  1. 1.
    task objects are randomly scattered on the ground (last two rows in Fig. 1),
  2. 2.
    distractor objects are present (last row in Fig. 1),
  3. 3.
    a human interrupts the task progress (in demo videos), or
  4. 4.
    objects to grasp have slightly different shapes (in demo videos).
We refer the reader to our demo videos in the supplementary file for more emergent generalization behaviors.
VIII-F Sim vs. Real Performance
Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variance in the RL training of the teacher policies. Additionally, because of the sim-to-real gap, real deployment can differ from simulation as well. To better understand how each technique impacts sim-to-real gap reduction, we showcase the simulation and real performance difference for all methods in Fig. 9. The sim success rate for each method denotes the average full task success rate over 100 episodes run in simulation, averaged across the three random seeds. The simulation environment follows the same randomization setting as training, with the policy rollout being fully greedy. The real success rate denotes the full task success rates from Table III.
Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasping, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up (without explicit rewards encouraging it), creating an exploration bottleneck for RL. Real-world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: the tip of the arm tends to shake more due to motor backlash, which are hard to be modeled precisely in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overloading. For No Perturbation, while one might expect perturbations to complicate RL learning, their inclusion results in much higher performance even in simulation. Without perturbations, the policy seems to converge to suboptimal solutions maybe due to a lack of exposure to a sufficiently large range of setups. Without visual augmentations (No Visual Aug), the simulation task is easier to be solved, and thus simulation success is artificially high, but real-world performance suffers when compared to the full method. RL training loss in the full SLIM method increases success in simulation (over Distillation Only) and achieves the highest success rate in real.
[…]
Figure 9: Full Task Sim and Real Success Rate Comparison. All success rates are the averaged success rate across three random seeds.
VIII-G Failure Modes
We also recorded failures for SLIM and report the most common cases here. In one case, the quadruped stopped a bit late during grasping and kicked the cube away. In another case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach. DropInto sometimes failed because of either early dropping right outside of the basket, or the quadruped kicking the basket away, right before dropping, when backing towards the basket. In rare cases where the gripper held only the top part of the cube, the gripper fingers completely blocked the cube from the camera view, and the robot got stuck hovering over the basket while holding the cube. See our failure demo videos for more details.
IX Limitations
Although not the main focus of this work, the current form of low-level policy introduces several limitations. One limitation is the applicable type of terrains. Currently the low-level policy is only trained on flat terrain in simulation. Although it shows the potential to work across different types of grounds (e.g., concrete, tile floor, carpet, mulch, etc.), it does not push the low-level policy learning to its full potential. We can potentially leverage the recent advances in locomotion policy training [67, 8] to further expand the ability and applicability of SLIM across more diverse types terrains, e.g., stairs, ramp, etc.
Another low-level related limitation is that the current form of low-level policy is not a wholebody policy. Although the current form of low-level policy is sufficient for the task considered in this work, by generalizing it to a wholebody policy [16], we can further enlarge its potential workspace.
As a pure sim-to-real approach that is trained from scratch, we did not leverage any pre-trained foundation models. As a result, we did not address a large amount of visual concepts and language variations and instead focused on the right amount of variations that are both practical and representative enough.
There are also a few limitations on the robot system side. Currently for the visual perception, we only used the RGB image from the wrist-mounted camera, which is the minimum setting that is sufficient for the current work. To further improve the robot’s perception and therefore the overall ability in terms of navigation and obstacle avoidance, it might be necessary to incorporate additional sensors for perception such as depth, stereo, lidar etc. In addition, this can also provide valuable information for the robot to improve its longer-term memories via SLAM-like mechanisms [63, 7].
X Conclusion, Discussion and Future Work
We present a complete legged mobile manipulation robot system for solving language-instructed long-horizon tasks in this work. The policy is trained via the sim-to-real paradigm, together with a progressive policy expansion-based teacher policy for solving the long-horizon task, followed with a distillation-guided RL approach for student visuomotor policy learning. We further identified and designed a suite of crucial techniques for reducing sim-to-real gap. Experimental results compared with a number of baseline methods verified the effectiveness of the proposed SLIM method. Moreover, real-world testing of SLIM across different scenes and spatial layouts shows that the proposed method can generalize well, further demonstrating the effectiveness of the whole system.
For future work, we would like to address the limitations discussed in Section IX, including further expanding the applicability of SLIM across more diverse scenarios by improving locomotion policy training [67, 8, 16]. Extending the visual and language diversity is another future work. Equipping the robot with additional sensors and enable it with SLAM-like ability [63, 7] is also an exciting future direction.
Appendix A RL Formulation
A-A Subtask Boundaries, Rewards and Behavior Priors
Subtask Boundaries.
Since we decompose the full long task into a sequence of subtasks, in practice, we need to decide the task boundaries in order to properly divide the full task. In the sequential decompositional case as described in Section VI, the task boundaries $\tau^{i}|\tau^{i+1}$ between two adjacent subtasks ($\tau^{i}\rightarrow\tau^{i+1}$) can be defined by the success condition $\mathcal{S}(\tau^{i})$ of the subtask $\tau^{i}$ (i.e., $\tau^{i}|\tau^{i+1}\equiv\mathcal{S}(\tau^{i})$), since the success of a subtask $\tau^{i}$ naturally leads to the subsequent subtask $\tau^{i+1}$.
We provide the success conditions of each subtask (therefore the subtask boundaries) in the sequel:
  1. 1.
    $\mathcal{S}$(Search): succeeds when the target object is in the view of the wrist-mounted RGB camera;
  2. 2.
    $\mathcal{S}$(MoveTo): succeeds when the target object enters a workspace of the robot within arm’s reach;
  3. 3.
    $\mathcal{S}$(Grasp): succeeds when the target object is in gripper and being lifted above the ground higher than 5CM and lasts at least two timesteps;
  4. 4.
    $\mathcal{S}$(SearchWithObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
  5. 5.
    $\mathcal{S}$(MoveToWithObj): succeeds when the target object is in gripper and the target basket enters a workspace of the robot within arm’s reach;
  6. 6.
    $\mathcal{S}$(MoveGripperToWithObj): succeeds when the target object is in gripper and the grippers moved to the top of the target basket;
  7. 7.
    $\mathcal{S}$(DropInto): succeeds when the target object enters the target basket.
Rewards.
We summarize the description of the subtask reward in Table IV. The task reward represents the sparse task success ($+1$) or failure ($-1$) reward. For timeout, we used a reward of 00. Apart from the sparse task reward, we also incorporated some shaping rewards, as shown in Table IV. Most of the shaping rewards are delta-distance based shaping reward, encouraging two position of interest to getting closer to each other. Arm retract reward is also an instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (please refer to Fig. 6 for a reference of the neutral arm pose). Intuitively, the arm retract reward will encourage the robot’s arm to rise up after grasping the object and move close to its neutral position. Alignment reward is also an instance is delta-distance based shaping reward, where the distance is defined based on the cosine between the left and right finger-tip-to-object vectors. Keep-grasping reward is a contact based reward issuing a penalty of $-0.1$ is the contact between gripper fingers and the object is lost.
TABLE IV: Rewards and Behavior Prior for Subtasks
Subtask Reward Behavior Prior
Search task rotational
MoveTo task + distance
Grasp task + distance + alignment stationary
SearchWithObj task + arm-retract rotational
MoveToWithObj task + distance + arm-retract + keep-grasping
MoveGripperToWithObj task + distance + keep-grasping stationary
DropInto task + distance stationary
Behavior Priors.
Another benefit of task decomposition is that we can easily incorporate different behavior priors to the policy. We used two types behavior priors in this work:
  • rotational: the forward velocity command ($\mathbf{c}[0]=0$) to the quadruped is kept as zero and therefore enable the quadruped to move in a rotational fashion. This is used as the behavior prior for search related subtasks (Search, SearchWithObj).
  • stationary: both the forward and angular velocity of the command to the quadruped are hardcoded as zero ($\mathbf{c}=0$). This encourages the robot to be stationary. It is used in manipulation related tasks (Grasp, MoveGripperToWithObj, DropInto).
In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g., the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior). The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution and being consistent with subtasks without these priors.
A-B Hierarchical RL and High-Level RL Training
As typical in standard hierarchical RL stuctures, the high-level policy and low-level policies are running at different frequencies. We refer one high-level inference step as a re-planning step. The high-level re-planning period is $5$, meaning the high-level policy (teacher or student) run once every $5$ low-level steps, and the same high-level output is used across this $5$ low-level steps re-planning period for the low-level policy.
Following this structure and given a frozen low-level policy, we are interested in training the high-level (teacher/student) policies. Because of this, the RL is essentially running at the high-level. Therefore, the per-(low-level-)step reward during each re-planning period is accumulated as the step reward for one high-level step, i.e.,
$r_{\rm high\_level\_step}=\sum_{i=1}^{5}r_{\rm low\_level\_step}^{i},$ (3)
and this accumulated reward $r_{\rm high\_level\_step}$ is used as the per-step reward signal for high-level RL training.
A-C Low-level Task and Reward
Low-level Tasks.
Details of low-level training tasks, commands, and corresponding rewards are summarized in Table V.
Quadraped Gait Control.
Motivated by [39], we also introduce gait alignment rewards that incentivize the robot to follow predefined gait patterns. The gait patterns are defined by a phase offset vector of length three, where each entry represents the phase difference of each of the three legs (excluding the first one) relative to the first leg (fore left). The range of each entry is $[-0.5,0.5]$, where 00 indicates a synchronous movement with the first leg and $-0.5$ or $0.5$ indicates a half cycle phase difference. The phase of a leg is evaluated independently for each leg, based on the leg’s schedule at current time. The schedule is a number in $[0,1]$. For gait alignment rewards design, schedule $1$ represents the “swing” phase where smaller movements in the X-Y plane will be penalized; schedule 00 represents the “contact” phase where lower contact force will be penalized; for any schedule between $(0,1)$, both aspects will be penalized with weights modulated by the schedule value.
For training of the gait following ability, we expand the input command $\mathbf{c}$ with randomly sampled phase offset vector $\mathbf{g}_{\rm offset}$ and gait cycle vector $\mathbf{g}_{\rm cycle}$. $\mathbf{g}_{\rm offset}$ is sampled from five predefined gait patterns, as summarized in Table VI. $\mathbf{g}_{\rm cycle}$ is a two-dimensional vector, with each dimension representing the frequency and the stance ratio respectively. The frequency takes values in $[2.0,4.0]$, indicating the number of cycles that each leg goes through per second. The stance ratio takes values in $(0.4,0.6)$, indicating the portion of time within a cycle when each foot is in contact with the floor.
After low-level training, we fix the gait pattern to be “trot”, i.e., $\mathbf{g}_{\rm offset}\!=\![-0.5,-0.5,0.0]$, and the gait cycle vector as $\mathbf{g}_{\rm cycle}\!=\![2.4,0.5]$. We have found that this gait is suitable for the mobile manipulation task tackled in this work in terms of movement efficiency and stability. For the stationary task of low-level training, we always set $\mathbf{g}_{\rm cycle}\!=\![0,1]$.
TABLE V: Low-level Training Tasks and Rewards
Task/Command Name Sample Range Reward
forward velocity $(-0.5,1.2)$ m/s $e^{-\frac{|v_{\text{forward}}-v_{\text{target}}|}{\max(v_{\text{target}},0.5)}}$
angular velocity $(-1.0,1.0)$ rad/s $e^{-|v_{\text{yaw}}-v_{\text{target}}|}$
stay stationary with probability $0.1$ $\begin{cases}-1&\text{if any foot lifted}\\ 0&\text{otherwise}\end{cases}$
balance and survival
tilt thresh: $0.6$
height thresh: $0.26$
$-(\mathbbm{1}_{\text{over-tilt}}+\mathbbm{1}_{\text{falling}})$
$-0.3\cdot\mathbbm{1}_{\text{killed}}$
TABLE VI: Predefined Gait Patterns
Gait Names Phase Offsets
pronk $(0.0,0.0,0.0)$
walk $(-0.5,0.25,-0.25)$
trot $(-0.5,-0.5,0.0)$
bound $(0.0,0.5,0.5)$
pace $(-0.5,0.0,-0.5)$
Low-Level Domain Randomization.
The domain randomization setting for low-level training is summarized in Table VII.
TABLE VII: Low-level Domain Randomization Parameters
Env Params Sample Range
base payload $(-0.5,3)$
torso mass center $(-0.15,0.15)$ for each dimension
feet friction $(0.25,1.75)$
feet softness (solimp[2]) [^3] Only the third entry of the solimp parameter of Mujoco is randomized. It is roughly proportional to the size of the deformation of its feet when the robot stands still on the ground. $(0.005,0.02)$
griper payload $(0,0.1)$
Kp of leg joints $(0.7,1.3)$
actuator gain of arm joints $(0.7,1.3)$
bumpiness of the ground $(0,0.15)$ for $90\%$, 00 for $10\%$
delay of sensors $(u-0.002,u+0.002),\ u\in(0.003,0.013)$
Appendix B Model Details
B-A Teacher (High Level)
The teacher network $\Pi\!=\!\{\pi^{k}\}_{k=1}^{K}$ is a set of sub-policies (referred to as policy set), as shown in Fig. 3. The effective size of the policy set is $K\!=\!8$ (including the Idle subtask).
Each policy net $\pi^{k}$ (for subtask $k$) is implemented as as a 3 layer MLP, with the layer size as $512$ ($[512,512,512]$). A squashed Normal projection layer (a layer for generating Normal distributions) is used as the output layer for each policy net following the standard practice [21], generating the action distribution conditioned on the current input.
The critic net for each subtask uses the same MLP structure and with the size of ($[512,512,512]$). Frame stacking is used with a stacking size of $5$.
The re-planning period is $5$, meaning that the high-level policy predicts once every $5$ low-level steps.
Figure A.10: The architecture of the student’s U-Net for predicting segmentation and depth maps.
Figure A.11: The architecture of the student’s policy input encoder that generates a compact observation embedding for RL.
B-B Student (High Level)
Maps Prediction U-Net.
An illustration of the U-Net [45] for predicting segmentation and depth maps is shown in Fig.A.10. The architecture details will presented according to input modality below.
  1. 1.
    Language. Due to the simplicity of our language instructions, we directly encode each instruction into a sequence of byte tokens. We set the max sequence length to be 100, padding shorter sequences with zeros while truncating longer sequences beyond this limit. Then each byte token is converted into an embedding vector of 128 dimensions by looking up a learnable embedding table. The sequence of language embeddings, with learnable positional encodings, is fed to an instruction encoder transformer [53] with self-attention. The transformer has 2 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 64 and a value dimensionality of 64. The MLP for generating the residual output has 128 hidden units. Finally, the output embeddings are average pooled to generate a compact instruction-level embedding of 128 dimensions
  2. 2.
    Proprioceptive state. We use a temporal stack of $N=5$. Different proprioceptive observations are simply flattened and concatenated together to produce a 1D robot state vector.
  3. 3.
    Vision. We use a temporal stack of $N=5$ RGB images for vision. The images are stacked along the channel dimension and fed to a U-Net encoder containing 6 convolutional blocks. Each block has a convolutional layer $(filters,kernel\_size,strides,padding)\triangleq(c,k,s,p)$, an instance norm layer [52], a ReLU activation, and a 2D max pooling layer (only starting from the second block). The convolutional layers are configured as $c=[32,64,64,128,128,128]$, and $k=3\times 3,s=1,p=1$ for all blocks. We use FiLM [42] conditioned on the instruction embedding to modulate each intermediate convolutional output, before it is fed to the next convolutional layer. After getting the output from the encoder, we concatenate it with the robot state vector and the instruction embedding, and feed it to the non-skip MLP. The MLP has two hidden layers of 512 units. Finally, the output from MLP is projected back to the output space of the encoder, and is fed to the U-Net decoder to produce the predicted map. The map has $M+1$ channels, where $M=4$ is the predefined maximal number of segmentation classes, and the remaining one channel is for depth prediction.
The maps prediction U-Net is trained by groundtruth segmentation and depth in simulation.
Policy Inputs Encoder.
An illustration of the policy input encoder network is shown in Fig. A.11. The network is responsible to digest predicted segmentation&depth maps and proprioceptive observations, and produce a compact representational embedding for downstream RL networks. Again, we use a temporal stack of $N=5$ for both segmentation and depth map inputs. After converting each segmentation map to a one-hot representation, we stack all maps along the channel dimension, resulting in an input of $5\times(4+1)=25$ channels. The input is fed to a ResNet [23] of 5 residual blocks, each block with 64 filters, a kernel size of $3\times 3$, and a stride of 2. The output from the ResNet is spatially flattened, resulting in a sequence of vision tokens. On the other hand, we use an MLP with one hidden layer of 256 units to project the concatenated proprioceptive observation into a token embedding of 256 dimensions. This proprioceptive token is appended to the vision token sequence which is fed to a representation transformer with self-attention. The transformer has 5 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 128 and a value dimensionality of 128. The MLP for generating the residual output has 256 hidden units. Finally, we take the transformed output corresponding to a special learnable CLS token as the final output of the entire policy input encoder.
Critic and Actor Networks.
Both the critic and the actor have three hidden layers of $(512,512,512)$. The actor outputs a squashed Gaussian distribution as in SAC [21]. Note that the RL and distillation gradients will be propagated all the way back to the policy input encoder for learning the representation.
Stationary Bit.
The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table IV). With this prior, the corresponding locomotion command distribution always has a zero mean and a small std. Taking argmax of the distribution, the teacher policy generates exactly zero locomotion command and thus the quadruped can remain stationary. However, if we directly distill this distribution into the student’s counterpart, it is almost impossible to guarantee a zero mode, due to the continuous nature of the distribution. Thus we employed a sationary bit trick to solve this issue. Specifically, we create an extra binary action distribution for the student for predicting whether the quadruped should be stationary or not, which is determined by checking if the locomotion command of the teacher is exactly zero. We then add a new prediction loss for learning this stationary bit, while masking out the KLD loss (only for the two locomotion command dimensions) if the stationary bit is true. For student inference, whenever the stationary bit distribution outputs 1, the locomotion command $\mathbf{c}$ will be overwritten to be zero.
B-C Low Level
For low-level training, we use PPO [47] with a regularized online adaptation module [16].
Policy Input and Locomotion Command.
As shown in Fig. 2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g., The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning. During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table V. The arm pose is also adjusted periodically within each episode, to make sure the low-level policy can follow locomotion commands while being robust against arm movements. For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm, empirically, we discover that it might cause the quadruped to shake more often, and decide to only use $\mathbf{s}_{\textrm{leg}}$. We also discover that the joint velocities in the hardware are delayed more than joint position readings from the motor sensors, possibly due to temporal smoothing. Since we do not know the exact algorithm that computes the joint velocities, we simply do not use them as input, and rely on feeding the neural networks the previous few frames for the policy to estimate the joint velocities.
After training, the low-level policy can be used to follow commands issued by an external command issuer. This can be either a higher level policy (as in SLIM or all the autonomous baselines), or human teleoperation (Human Teleop) in this work.
Actor and Value Networks.
Both the actor and value networks have three hidden layers of size $(256,256,256)$. The actor network outputs a Beta distribution instead of the typical Gaussian. For numerical stability, we set the minimum value of the shape parameters of the Beta distribution to $1$.
Online Adaptation Module for Sim-to-Real Transfer
. Similar as in [16, Fig.2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder. The privileged encoder takes privileged information as input and predicts a latent representation of environment extrinsics, while the adaptation encoder predicts the same latent representation depending only on robot sensory observations. Both encoders are trained to match each other and jointly with the RL objective. The privileged encoder network has two hidden layers of $(128,64)$ and the adaptation encoder has hidden layers of $(256,128)$.
B-D Model Parameter Count
We summarize in Table VIII the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
TABLE VIII: Number of Trainable Parameters.
Component Approx. Param. Count
Low Level $3$M
Teacher High Level (all subtasks) $10$M
Student High Level $16$M
Our deployed policy (“Low Level” + “Student High Level”) has a total parameter count of less than 19M.
Appendix C Important Training Hyperparameters
Below we list some important training hyperparameters of SLIM.
Teacher High Level.
We used four 3090 GPUs for training, with a batch size of $2000$, a learning rate of $1\times 10^{-4}$, gradient clipping of $0.1$ with the Adam optimizer, $120$ parallel environments, and training iterations of $4$M (corresponding to about $60$M environment steps). The total training wall-clock time was about 1 week.
Student High Level.
We used 60 parallel environments to collect rollout data for training the student, where the data was stored in to a replay buffer with a capacity of $600$k environment steps. The full policy was trained on four 3090 GPUs for about 2 weeks, for a total of $1.3$M iterations. This amounts to roughly $40$M environment steps in total. We used a batch size of $64$, a learning rate of $1\times 10^{-4}$, gradient clipping of $100$, and the Adam optimizer. For the mixed rollout schedule, we kept the probability $\beta$ of sampling from the student policy as $0.5$ until the $750$K-th iteration, and linearly increased $\beta$ so that it reaches $1.0$ at the $1.25$M-th iteration. The distillation loss weight $\alpha$ in Equation 2 was set to $0.01$. The fixed modal dispersion $\sigma$ for distilling the teacher policy was $0.05$.
Low level.
We used 5000 parallel environments to roll out and train the low-level policy. The policy was trained on one 3090 GPU for about three days for a total of 4B environment steps. We used a batch size of 5000. The Adam optimizer had an initial learning rate of $2\times 10^{-4}$ and a step decay to its $1/5$ after $70\%$ of the total training progress.
Appendix D Simulator Choice
All simulations were done in the MuJoCo simulator [51]. At the beginning of the project, we evaluated several popular simulators at the time, including MuJoCo [^4] https://mujoco.org/, Isaac Gym [^5] https://developer.nvidia.com/isaac-gym, and PyBullet [^6] https://pybullet.org/wordpress/. We settled on MuJoCo due to its reasonably fast simulation (on CPU) and rendering (on GPU) at the same time.
Besides camera observation rendering for student training, due to partial observability, we also rely on segmentation rendering to check which objects are in the camera view and use it as privileged information for the teacher, during both teacher and student training. Thus RGB and segmentation rendering efficiencies are critical to our simulator choice.
Appendix E Sim-to-Real Techniques
E-A Dynamics Gap Reduction Techniques
The dynamics sim-to-real gap is the mismatch between the transition function $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ of the simulation versus that of the real world. It can be due to the mismatch in the physical properties of objects and motors, or by things that are not properly simulated, such as friction and backlash. This difference in dynamics can cause policies trained purely in simulation to fail when deployed in the real world, especially if the task requires accurate motor control, such as grasping a small object. We use the following key techniques to address the dynamics gap.
Arm PID Control to Minimize Tracking Errors.
An arm joint tracking error $\mathbf{e}_{\textrm{arm}}$ is the positional difference between the arm command of the previous timestep and the joint positions of the next timestep,
$\mathbf{e}^{t}_{\textrm{arm}}=\mathbf{s}^{t}_{\textrm{arm}}-\mathbf{a}^{t-1}_{\textrm{arm}}.$
Prior works [16] have used PD control for the arm joints. PD control needs to have large tracking errors, the difference between the target and the current joint positions, to compensate for lifting the arm and any grasped object against gravity.
Intuitively, we know that if the tracking error $\mathbf{e}_{\textrm{arm}}$ is equivalent between sim and real for all visited states, then there is essentially no dynamics gap. With this, one solution would be to match the tracking errors as closely as possible through accurate embodiment modeling with low-level PD control tuning. In practice, achieving this alignment for the entire state distribution requires nontrivial work, given the nonlinearity and drifting of motor dynamics due to effects such as motor backlash, wear-and-tear, etc.
Rather than matching non-zero tracking errors, another approach is to simply achieve near-zero tracking errors in both sim and real. In other words, whatever arm command is outputted, the resulting arm position should be roughly achieved $\mathbf{s}^{t}_{\textrm{arm}}\approx\mathbf{a}^{t-1}_{\textrm{arm}}$. In fact, prior studies have shown that tracking error minimization is one of the most reliable strategies for enabling successful sim-to-real transfer [3]. To do so, we use a PID joint position controller for our manipulator in both simulation and in real, with a small joint position change limit $z$ (Section III-A) to ensure minimal tracking error for a wide variety of pose configurations. Our tuned PID controller executes smoother and slower trajectories than those from the PD controller, allowing safe and accurate deployment.
We stay with PD control for the quadruped motors as the hardware driver does not support PID.
Arm Control Perturbation.
We add ±$\pm 0.02$ noise to the arm joint position targets uniformly randomly, sampled every 50 high-level control steps or 5 seconds real time. This noise is applied to both teacher and student training. The noise level is increased to ±$\pm 0.05$ after 3M iterations for teacher training. This is to make the learned policy to be robust w.r.t. different sources of control noise arising when transferred to real.
Arm Mount Perturbation.
We randomly perturb the arm mount position and yaw at the beginning of each episode to make the learned policy robust to sim-to-real gaps that may arise from the actual mounting accuracy, torso height variations, etc. In this work, we sample perturbations for $x$, $y$, $z$ and yaw $\theta$ from the following ranges: $x:\pm 1$cm, $y:\pm 1$cm, $z:\pm 1$cm, $\theta:\pm 2\times 10^{-3}$ rad.
Object Perturbations.
Due to the relatively deterministic nature of simulations (in contrast to the high-entropy real world), we observed that policies often converge to deterministic strategies. For example, for picking up an object, a policy may learn to always position itself so that the object has the same relative position to its base. This deterministic “memorizing” behavior can result in task failure if the conditions for success are even slightly off. Therefore, we enforce learning robust, reactive policies by perturbing objects of interest during various stages of task execution. Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robustness of the policy (as the state coverage is enlarged).
We perturb the object within circle with a radius of $r_{\rm perturb}$ centered around the current position of the object. We randomly sample a position within this circle and check its eligibility. If it touches or within the space of robot torso projected on to the ground plane, then this perturbation position is regarded as invalid. Otherwise it is regarded as valid. If a valid perturbation is found within 10 trails, then we transport the object to the new location as a form of object perturbation. Otherwise, we keep the position of the object unchanged. In this work, $r_{\rm perturb}\!=\!10$cm is used.
Stationary Manipulation.
We make the quadruped stand as still as possible when manipulating objects, i.e., during grasp or drop into. This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed. To achieve this, when training low-level policies to follow locomotion commands, we also encourage the quadruped to enter a stationary mode with a balanced stance when the speed command is close to zero. Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers, arm motors and the surrounding.
Reducing Quadruped Shaking
When transitioning from locomotion to stationary, the quadruped after adding the arm on top can shake, sometimes quite violently, before settling. This only happens in real deployment, not in simulation. Our experiments seem to show that hiding the arm command and observations from the low-level locomotion policy during training and deployment reduces most of the shaking. Shaking when transitioning to stationary is still present in most low-level policies. For the experiments in this paper, we ran low level training for ten random seeds, and chose three of the best ones (less shaking and more stable stationary pose) to continue high level policy training and evaluation.
Dynamics Sim-to-Real Gaps Not Addressed.
Unlike prior work, we did not explicitly model the motor dynamics, and thus the effects of backlash and motor operating near torque limit boundaries are not modeled. For the leg motors of Go1, we follow existing work which uses P control plus damping to mimic the physical PD controller. Damping is applied on the absolute joint velocity of the motor, while the D factor in PD control is applied on the difference between the targeted and the actual joint velocities. During fast leg joint movements, this difference can cause a large sim-to-real gap.
E-B Vision Gap Reduction Techniques
The second source of sim-to-real gap for image-conditioned policies is the vision gap. This gap is caused by a visual distribution between simulated and real world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios. Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
Image Domain Randomization.
Inspired by SECANT [12], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim-to-real transfer. Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image. We choose not to introduce higher-level image randomization/generation [60] because it might impact image semantics.
Random Textures.
We spawn the robot on a flat ground plane and randomize the ground texture. The texture can be arbitrary and generally has no specific semantics, as long as they provide enough visual distraction to the student. This distraction will force the student to learn to focus on the most important visual features that are related to the task at hand: objects with reasonable sizes and target colors. When deployed in real, we hope that this visual focus will get transferred and help the robot locate objects.
Random Spatial Augmentation.
To make the visual representation module robust, we incorporate spatial augmentations in the visual representation learning [57, 58].
Random Background Objects.
To further increase the visual complexity of the surrounding environment, we randomly sample and spawn a set of background objects in the room around the robot’s initial position (Figure A.12). There are two categories of background objects: 1) primitive shapes (cuboid, cylinder, and sphere) and 2) Google scanned objects [10]. We make sure that these objects only serve the purpose of distraction and the student will not confuse them with task objects. It could collide with an background object during training with some penalty reward. When spawning an object, we also randomize its size, orientation, and texture (only for primitive shapes).
Refer to caption Refer to caption
Figure A.12: Example Training Scene with Randomized Objects. Left: isometric view; Right: birds-eye view.
No Camera Depth Data.
For our setting, language-guided visual recognition in the RGB space is the key to finish the task. Without pretrained vision models, this sim-to-real transfer in RGB has to be supported by our method. Given this, we choose not to use camera depth data because this will introduce an extra sim-to-real gap, which has to be reduced with denoising or hole-filling techniques [25]. We observe that depth can be reasonably estimated from stacked RGB images, even with monocular vision.
Color Modeling.
Vision sim-to-real, especially for colors, is notoriously difficult [44]. We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow. These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training. To obtain the randomization range, we randomly sample 1000 pixels for each color, as shown in Fig. A.13. Afterwards, we compute the means and standard deviations of the samples. HSV values are then sampled during training from a uniform distribution with a matching mean and $\pm$ bounds that cover the real-world distribution.
Figure A.13: Color samples. Each color had 1000 randomly sampled pixels used to estimate the distribution in simulation.
Appendix F Robot Implementation Details
F-A Control Latency Management
We measured the latency along the full control cycle, including motor sensor reading ($\sim$3 ms), camera observation (30$\sim$40 ms), network connection ($\sim$0.2 ms with cable and $\sim$5 ms on Wi-Fi), and model inference (20$\sim$40 ms for the high level  [^7] Later with TensorRT, we were able to reduce this latency to around 10 ms. However, the experiments in this paper were conducted without this option back then., and $\sim$5 ms for the low level). We implement the following methods to address such delays as much as possible, while properly modeling them in our simulator.
  1. 1.
    Asynchronous low-high inference: most of the model latency comes from the high-level inference, which takes around 20 ms, but can sometimes peak to 40 ms. This could delay the low-level policy (50 HZ) considerably and make the robot falter, if the two levels are executed in sequence. Alternatively, in the asynchronous mode, the low-level policy will reuse the previous locomotion command if the new one has not arrived yet, and thus still maintain its 50 HZ frequency. This leads to a more stable gait of the quadruped.
  2. 2.
    Ethernet cable connection to the robot: Wi-Fi connection is more convenient than cable. However, we find the Wi-Fi connection between Go1 and the laptop to be unstable, sometimes spiking to a latency of 100 ms, which is unacceptable for low-level motor control. It is possible to have a Wi-Fi solution with more powerful and stable hardware, or to run low-level inference on Go1 instead of over Wi-Fi, but we simply rely on cable connection in this work.
  3. 3.
    Sensor delay randomizations: motor sensor reading delay is sampled from $(3,13)$ ms per episode with $\pm$ 2ms variation uniformly per control step. Camera observation delay is sampled from $(30,50)$ ms uniformly per step with observations taken at 60 HZ. Low-level control delay is sampled from $(5,7)$ ms uniformly per step.
  4. 4.
    High-level inference delay: we trained policies without high-level inference delay turned on, and compared the performance under different extra test-time high-level inference delays. We did not find clear performance differences in terms of the success rate. Therefore, we did not use high-level control delay in policy training in this work. Meanwhile, it is an option that is provided in our implementation and can be turned on whenever necessary.
Figure A.14: Asynchronous System Architecture. Note that “Onboard*” refers to the low-level controller frequency being unknown due to the use of closed-source software. Furthermore, when using the WidowX’s PID controller, we also take advantage of Dynamixel’s onboard profiling for smooth control.
F-B Safety Measures for the Physical Robot
To reduce damages to surroundings and the robot itself, we employed the following safety measures.
  1. 1.
    We first addressed all sim-to-real gaps of the arm alone by training and deploying a tabletop policy. Similarly, we trained and deployed a standalone quadruped policy, without the mounted arm. Only after the two standalone policies were transferred well, we started training and deploying the whole-body policy. This helped us avoid potential catastrophic failures from the first-time deployment of a complex quadruped manipulation system.
  2. 2.
    We used PID control for the arm (Section E-A), with a reasonably small action clipping, resulting in much smoother and more reliable arm movement. This strategy reduced damages to the arm motor gears significantly.
  3. 3.
    The randomizations of motor strength, payload, center of mass, arm mounting position, etc. (Section E-A), though making simulation training more difficult, actually allowed the policy to adapt to the real world more easily, resulting in very accurate grasping executions. Thus the arm did not push into the ground or the target objects due to dynamics mismatch, avoiding motor damages.
  4. 4.
    As high-level inference can sometimes take over 40ms to finish, we adopted asynchronous low-high inference (Fig. A.14) to make sure that the high-level inference did not prevent the robot from receiving closed-loop low-level commands. Otherwise, the quadruped could topple over and damage components mounted on top.
  5. 5.
    We also enabled deployment termination with a single button press. When latency spikes, especially when the deployment machine suddenly freezes or the network disconnects, our driver on Go1 will automatically stop the quadruped. This preventative stop was done using a scripted policy and it prevented damages to the hardware in most cases.
  6. 6.
    To prevent the robot from wandering off and bumping into surroundings, during training we incorporated a rotational search behavior as explained in Section A-A. This was done for all baseline methods and SLIM.
  7. 7.
    We attached a leash to the quadruped to prevent it falling down due to any reason. Later when the policy became more reliable in real, we removed the leash.
  8. 8.
    Our 3D-printed camera mount (Fig. 6) was designed to break easily, sacrificing itself to protect the camera in the event of a side fall.
  9. 9.
    During deployment, we frequently monitored the temperature of Go1’s motors and took a break if necessary. A very hot Go1 motor will result in unpredictable, dangerous locomotion behaviors.
Appendix G Real-World Experiment Details
G-A Containers and Graspable Objects
We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Fig. A.15. The size of the basket is about $16{\rm cm}\!\times\!16{\rm cm}\!\times\!16{\rm cm}$. The size of the cube is about $2.5{\rm cm}\!\times\!2.5{\rm cm}\!\times\!2.5{\rm cm}$ and the weight of the cube is about $9{\rm g}$.
[…]
Figure A.15: Containers and Graspable Objects.
G-B Real-world Evaluation Protocols
We detail the protocols used for real-world quantitative evaluation in the main paper (Table III).
TABLE IX: Scene Setups for Real-World Quantitative Evaluations. Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$. Uppercase letters denote baskets and lower case letters denote the objects to be grasped. The color of an object is denoted both by the background color and the character (e.g., g denotes the green cube and R denotes the red basket.)
Scene setup Instruction: ”Drop x into Y” (x $\mapsto$ Y)
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
b $\mapsto$ R
b $\mapsto$ B
[…]
y $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
  1. 1.
    For each method, we run 20 episodes with different initial scene states;
  2. 2.
    We apply a time limit of 90 seconds for each evaluation episode.
  3. 3.
    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Fig. 7, as it is easy to ensure accuracy of the scene setup in terms of spatial positions. In practice, we mark the positions for the robot at the center of the square. Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets. The size of the square is about $2\times 2$ meters.
  4. 4.
    We follow a scene initialization procedure as detailed in Table IX across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
Appendix H Visualizations
[…]
Figure A.16: Policy Trajectory Visualization. For visualization purpose, we removed background objects. We also visualize the EE-trajectories. Trajectories corresponding to different subtasks are rendered with different colors. The ego-centric camera view is shown on the top-right of each image. The red colored arrows denote the direction of movement.
[…]
Figure A.17: Maps prediction results for some key frames of a deployment session. Time flows from top to bottom, and each row corresponds to one time step. From left to right, the columns represent the RGB frame, the predicted depth map, the predicted segmentation mask, and our added note for better interpretation of different stages.
H-A Visualization of Policy Behaviors
The visualization of the policy behavior in real is shown in Fig. 1. We also provide demo videos in the supplementary file. Here we further visualize the policy behavior across the complete long-horizon task with the end-effector (EE) trajectory being rendered. To do this, we run the policy in the simulator, and render the images from a fixed top-down perspective and overlay the EE-trajectory. The EE-trajectory is color-coded in such a way that the segment corresponding to different subtask is rendered with a different color. The behavior of the policy is shown in Fig. A.16. To make the visualization clear we did not add background objects in the simulation for this episode. The robot start with a randomized position and yaw within a region (Start phase). A red cube and a green cube as well as a red basket and and blue basket are used. Their positions are randomized around the Standard layout and their yaw angles are also randomized. Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube). Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f. Fig. A.16-Search). Once the target object is in view, the robot will move towards it (Fig. A.16-MoveTo). Once the robot moves close to the target object, it will start to grasp (Fig. A.16-Grasp). During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it. After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper. As shown in Fig. A.16-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory). At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE-trajectory in Fig. A.16-SearchWithObj. Once the target basket is in view, the robot will move towards it (Fig. A.16-MoveToWithObj) and drop the cube in gripper into that basket (Fig. A.16-DropInto).
H-B Visualization of Predicted Segmentation and Depth Maps
In Fig. A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session. We emphasize again that the robot was deployed in this zero-shot scenario never modeled by our simulator, and no model finetuning was performed. Despite this, we can see that the prediction results are high-quality, given only a very short history ($0.5$ seconds) of a single wrist-mounted RGB camera and the robot proprioceptive state. While searching for the target cube, the perception model had to locate it with just a handful of pixels (first row). Right after the gripper holding and lifting the cube above the ground, the model correctly predicted that the target had switched from the cube to a basket (cube’s green color turned red; third row). Moreover, while searching for the target basket in the outdoor environment, the cluttered background did not confuse the depth and segmentation prediction (fourth row). The perception model was also able to correctly predict the segmentation mask for the robot itself (gripper fingers and quadruped in orange; fifth row). Overall, we find that in practice, this high-quality visual information bottleneck not only greatly improves the generalization ability of our RL policy, but also facilitates our debugging of the entire robotic system as it separates the errors of the perception model from those of the policy.
Appendix I Human Teleoperation
We list failure modes of the Human Teleop baseline to give a more complete picture of how difficult our task is, and to show that SLIM’s performance is non-trivial.
  1. 1.
    The most common failure mode for teleoperation is task timeout. This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time for searching and grasping again. Failing to grasp fully is in turn due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision.
  2. 2.
    The arm can have a twisted pose due to IK near singularities. This pose either overloads the arm motors, or if not, flips the camera upside down, making the operator difficult to perceive the environment.
Overall, human teleoperation is less accurate and slower than an RL policy, due to the low-level policy’s shakiness, the difficulty of simultaneously moving the end effector during locomotion, and sometimes human errors resulting from lost focus after a long time. Besides being less time efficient, human teleoperation seems to be less energy efficient than our RL policies as well. While the RL trained high-level policy can run smoothly on a particular low-level policy, human teleoperation on the same low-level policy sometimes triggers over-current protection of Go1. This often happens when an over-extended arm is raised, causing Go1 to lose control of its leg motors. In evaluation, we ignored this Go1 hardware failure, and let the human operator simply try from start again.
Appendix J Long-Horizon Task Learning
J-A Long-Horizon Task with Bottleneck States
The challenges of long-horizon task for exploration and learning have been discussed in Section IV. Here we want to further highlight that the existence of bottleneck states, which is orthogonal to long horizon, further exacerbates the difficulties.
[…]
Figure A.18: Long-Horizon Task with Bottleneck States. Left: Top-down view of the end effector trajectories. Right: The same set of trajectories in 3D space.
A bottleneck state is a milestone that must be accomplished before reaching the final task success. Therefore, given two tasks of the same horizon length, the one with the presence of bottleneck states will typically be more difficult than the one without.
To visualizing the bottleneck states for the tasks considered in this work, we use spatial proximity of a key position (end effector/EE position) as a proxy of state proximity. Therefore, positions that are spatially close to each other can be roughly interpreted as with similar states in terms of accomplishing the mobile manipulation task under consideration. Under this setting, for the same task, if there is a common segment along the way with a compact spatial support, it can be regarded as a bottleneck state.
We visualize the trajectories from three episodes in Fig. A.18, given a fixed object layout and different starting positions. As we can see from the figure, although with quite different starting positions, all the trajectories converge and intersect at the same location marked by b in the figure, corresponding to the state for grasping the target object.
To achieve the eventual full task success (drop the target object into basket), a successful grasping of the target object has to be accomplished as a prerequisite. Given that grasping small objects requires high precision control (as shown by the inserts in the left part of Fig. A.18), this forms a clear bottleneck state, increasing the difficulties of exploration and learning, especially for the part of state space that is only available after passing the bottleneck state.
The existence of bottleneck states coupled with long task horizon pose a great challenges to standard algorithm, as shown by the results in the subsequent section.
Figure A.19: Standard Method v.s. Progressive PEX. We compare the learning behaviors of Standard method (without progressive policy expansion) and the Progressive PEX approach in terms of the task progress as training progresses.
J-B The Effectiveness of Progressive Policy Expansion
Section IV introduces the Progressive Policy Expansion approach (Progressive PEX) for long-horizon task learning. The comparison of Progressive PEX and the approach using a standard network without progressive policy expansion (Standard) on solving the long-horizon task are shown in Fig. A.19. We show each policy’s ability in solving the long-horizon task throughout the training process, measured by the number of subtasks that can be solved ($y$-axis, Long Horizon Task Progress). We order the subtasks following their respective orders appeared in the full task. We show the mean and standard deviation calculated across 3 seeds.
As can be observed from the figure, Standard approach starts with a similar trend in learning to solve subtasks as Progressive PEX. However, after learning to solve a few tasks, Standard appears to be saturated in terms of the learning progress, likely due to the inability of continual exploration and loss of plasticity [35, 36].
In contrast, Progressive PEX can keep pushing the frontier on the long-horizon task progress, learning to solve the subsequent new subtask that appears along the way towards solving the full task.